### v2

# Structural Cointegration of the Paleoclimate

Abstract
This paper investigates the coupled dynamics of temperature and atmospheric CO2 concentrations during Pleistocene glacial cycles using a Structural Vector Error Correction Model (SVECM). To address the controversy regarding causal leads and lags, we develop a structural identification framework that utilizes Milankovitch orbital parameters as instrumental variables to disentangle the Instantaneous Greenhouse Effect (IGE) and the Instantaneous Carbon Feedback (ICF) at a millennial resolution. Focusing on the gap-free interval of Termination II, we employ a Causality Allocation Curve (CAC) derived from system-wide residuals to identify the causal structure that maximizes statistical consistency with the observed record. This approach allows for the estimation of Earth System Sensitivity (ESS) based purely on observational data and the decomposition of warming dynamics into structural momentum and equilibrium adjustment. By demonstrating the consistency of these paleo-derived structural parameters with modern instrumental records, this study establishes a robust econometric foundation for integrating long-term climate analysis with contemporary climate monitoring.
Introduction
The strong coupling between atmospheric CO2 concentrations and global temperatures over the glacial-interglacial cycles of the Quaternary period is one of the most striking features of the Earth’s climate system. High-resolution ice-core records from Antarctica, spanning the last 800,000 years, have provided unequivocal evidence that CO2 and temperature covary with remarkable precision across multiple terminations and glaciations (Jouzel et al., 2007; Lüthi et al., 2008). This tight correlation forms the empirical basis for our understanding of the greenhouse effect as a fundamental regulator of the Earth’s long-term climatic state.
However, the precise causal architecture of this relationship remains a subject of intense debate in paleoclimatology. Early investigations into the timing of these transitions suggested that Antarctic temperature increases preceded CO2 rises by several centuries during glacial terminations, leading to the hypothesis that CO2 acts primarily as an amplifier of orbitally-induced warming rather than its primary driver (Caillon et al., 2003). Conversely, recent re-evaluations of ice-core gas-age chronologies have argued that the perceived lag may be a statistical artifact of the firn densification process, suggesting that CO2 and temperature are effectively synchronized at the millennial scale (Parrenin et al., 2013). This “lead-lag” controversy is not merely a chronological puzzle; it strikes at the heart of our ability to quantify the true climate sensitivity and the structural feedbacks inherent in the Earth system.
The fundamental difficulty in resolving this causality lies in the simultaneous nature of the climate-carbon coupling. Within the observational resolution of the paleoclimatic record, temperature and CO2 operate in a continuous feedback loop: warming triggers the release of CO2 from the oceans and terrestrial biosphere, while increased CO2 further warms the planet through radiative forcing. In econometric terms, this constitutes a problem of simultaneity and endogeneity, where standard correlation analyses or simple Granger-causality tests are insufficient to isolate the structural parameters of the system. Without a robust identification strategy, it is impossible to determine whether the observed covariance is driven by an instantaneous greenhouse effect or by a temperature-driven carbon response.
In this paper, we address this identification challenge by applying a Structural Vector Error Correction Model (SVECM) to the Pleistocene ice-core record. Our approach builds upon the methodology of structural macroeconometrics, utilizing the strict exogeneity of Earth’s orbital parameters—Milankovitch forcing—to achieve causal identification. We distinguish between two contemporaneous mechanisms operating within a 1,000-year time step: the Instantaneous Greenhouse Effect (IGE), representing the immediate thermal response to CO2 innovations, and the Instantaneous Carbon Feedback (ICF), representing the carbon cycle’s response to warming.
By analyzing the gap-free record of Termination II, we demonstrate that the Pleistocene climate is characterized by a null Instantaneous Greenhouse Effect ($a_{12}\approx 0$) at the millennial resolution. Instead, we show that the greenhouse effect is structurally localized as a powerful delayed feedback. Our model identifies a structural lag of 1,000 years, after which a CO2 doubling triggers a cumulative temperature increase of approximately 11.8 K, comprising a 10.5 K structural momentum and a 1.3 K long-run equilibrium adjustment.
Crucially, we validate these paleo-derived structural parameters by projecting them onto the modern era. We show that the centennial warming rate derived from our Pleistocene model accurately recovers the observed instrumental warming of the last 150 years. This consistency suggests that the fundamental coupling between the thermal and carbon cycles has remained stable across vastly different climatic regimes. By providing a definitive empirical verification of the structural lag in the greenhouse effect, this study offers a new framework for reconciling the millennial-scale dynamics observed in the ice cores with the centennial-scale warming observed in the modern anthropogenic era.
While earlier studies, such as Kaufmann et al. (2006), have successfully applied VECM to instrumental climate records to identify long-run relationships, they relied on pre-defined radiative forcing formulas and reduced-form estimations. By contrast, our study extends this econometric framework into the structural domain (SVECM). By utilizing the orbital forcing as a source of exogenous variation, we identify the distinct structural parameters of the Earth system—IGE, ICF, and the structural lag effects. This allows us to estimate an Earth System Sensitivity (ESS) that is derived purely from observational data, consistent with the long-term feedback framework proposed by the PALAEOSENS project (Rohling et al., 2012).
Model
In this section, we delineate the mathematical and physical foundations of our empirical approach. Given the high degree of persistence and the inherent feedbacks between temperature and greenhouse gas concentrations observed in ice-core records, a robust identification strategy must account for both long-term equilibrium states and short-term structural innovations. [^1] For a more exhaustive mathematical background on the Vector Autoregressive (VAR) and Vector Error Correction (VEC) frameworks, including the properties of non-stationary systems and the derivation of the long-run permanent component ($\Xi$), the reader is referred to Appendix.
Vector Error Correction Model
A standard Vector Autoregressive (VAR) model typically assumes that the underlying time series are stationary. However, paleoclimatic proxies such as temperature and CO2 concentrations exhibit the properties of non-stationary, $I(1)$ processes. Estimating a VAR model in levels for such variables carries the risk of spurious regression, where the high degree of persistence leads to statistically significant correlations that do not represent true physical causalities.
To circumvent this risk while preserving the critical information contained in the levels of the variables, we employ a Vector Error Correction Model (VECM). This framework is appropriate when variables are cointegrated, meaning that while individual series may drift non-stationarily, a specific linear combination of them remains stationary. This stationary combination represents a long-run equilibrium or a physical bond that has persisted through numerous glacial-interglacial cycles. The VECM is formulated as follows:
$\Delta y_{t}=\alpha\beta^{\prime}y_{t-1}+\Gamma_{1}\Delta y_{t-1}+\dots+\Gamma_{p-1}\Delta y_{t-p+1}+\epsilon_{t}$ (1)
where $y_{t}=(T_{t},\ln C_{t},M_{t})^{\prime}$ is the vector of endogenous variables representing temperature ($T_{t}$), log-transformed CO2 concentration ($\ln C_{t}$), and orbital forcing ($M_{t}$). $\Delta$ denotes the first-difference operator, and $\epsilon_{t}$ is a vector of reduced-form residuals.
The matrix $\Pi=\alpha\beta^{\prime}$ contains the essential information regarding the cointegrating space. The vector $\beta$ defines the long-run equilibrium toward which the climate system gravitates over centennial to millennial scales. The existence of a stable $\beta$ suggests a fundamental consistency in the coupling of the thermal and carbon cycles across diverse climatic regimes.
The matrix $\alpha$ represents the speed of adjustment or the restoring force of the system. When a stochastic shock—such as an orbital perturbation—displaces the system from its equilibrium path, the error correction term ($\beta^{\prime}y_{t-1}$) becomes non-zero. The coefficient $\alpha$ dictates how rapidly the system initiates corrective dynamics. This mechanism constitutes the mathematical representation of the climate system’s self-regulation.
Structural Identification and Causality Allocation
To isolate structural causalities from the reduced-form correlations, we assume that the VECM is derived from an underlying Structural VAR (SVAR) model. This approach follows the framework established by Sims (1980), which provides a basis for analyzing endogeneity and structural innovations within a system of interrelated time series by utilizing the residuals of the estimated model. By applying this structural methodology to the paleoclimatic system, we can define a model that explicitly accounts for the contemporaneous interactions between temperature and carbon cycles:
$(I-A_{0})\Delta y_{t}=(I-A_{0})\alpha\beta^{\prime}y_{t-1}+\Gamma^{*}1\Delta y{t-1}+\dots+u_{t},$ $u_{t}\sim\text{IID}(0,I)$ (2)
The matrix $A_{0}$ represents the simultaneous feedbacks within the same 1,000-year time step. The reduced-form residuals $\epsilon_{t}$ are related to the structural shocks $u_{t}$ via the relation:
$u_{t}=(I-A_{0})\epsilon_{t}$ (3)
Here, the elements $a_{ij}$ of the structural matrix $A=I-A_{0}$ designate the contemporaneous causal effects.
Following physical constraints, we assume that temperature and CO22 do not affect orbital forcing ($a{31}=a_{32}=0$), and that orbital forcing does not directly affect CO22 concentration (a23=0a{23}=0). Under these restrictions, Equation (3) yields the following system of structural innovations:
$\begin{aligned} u_{1t}&=\epsilon_{1t}-a_{12}\epsilon_{2t}-a_{13}\epsilon_{3t}\\ u_{2t}&=-a_{21}\epsilon_{1t}+\epsilon_{2t}\\ u_{3t}&=\epsilon_{3t}\end{aligned}$ (4)
where $a_{12}$ denotes the Instantaneous Greenhouse Effect (IGE)—the immediate temperature response to a change in CO22 within the same millennium—and a21a{21} denotes the Instantaneous Carbon Feedback (ICF). Importantly, $a_{12}=0$ corresponds to a null-IGE identification, where CO2 increases exert no additional warming effect at the contemporaneous resolution.
Identification of these parameters requires the orthogonality of structural shocks ($\mathbb{E}[u_{it}u_{jt}]=0$ for $i\neq j$). Substituting Equation (4) into the condition $\mathbb{E}[u_{1t}u_{2t}]=0$, we derive the Causality Allocation Curve (CAC):
$(1+a_{12}a_{21})s_{12}-a_{21}s_{11}-a_{12}s_{22}=0$ (5)
This hyperbola represents all causal allocations $(a_{12},a_{21})$ consistent with the observed residual covariances $s_{ij}$.
Impulse Response Analysis
To evaluate the dynamic stability and long-term consequences of structural shocks, we employ impulse response functions (IRFs). By utilizing the moving average representation of the VECM, the level of the variables at horizon $h$ is expressed as:
$y_{t+h}=\Psi_{0}u_{t+h}+\Psi_{1}u_{t+h-1}+\dots+\Psi_{h}u_{t}$ (6)
The matrix $\Psi_{j}$ represents how a structural innovation at time $t$ propagates through the simultaneous interactions and the long-run adjustment mechanism. We simulate two hallmark scenarios to quantify climate sensitivity:
  1. 1.
    A CO2 doubling shock: Represented as a structural innovation in the carbon cycle ($u_{2t}$). In our logarithmic specification, a doubling is equivalent to an instantaneous shock of magnitude $\ln(2)\approx 0.693$. This allows us to trace the temperature response via the IGE and subsequent feedback loops.
  2. 2.
    A 2K temperature shock: Represented as an abrupt warming ($u_{1t}=2.0$). This scenario quantifies the magnitude of the resulting CO2 release, identifying the strength of the ICF and the efficiency of the error correction process.
Our VECM is estimated on a 1,000-year grid to ensure a high signal-to-noise ratio in the ice-core proxies. While the structural dynamics are defined on this millennial scale, we utilize spline interpolation when presenting the IRFs to facilitate comparison with higher-frequency records without altering the fundamental parameters.
As derived in Appendix A, the presence of cointegration ensures that temporary structural shocks lead to a permanent shift in the climatic baseline. The long-run impact matrix, defined as $\Xi=\lim_{h\to\infty}\Psi_{h}$, represents the total accumulated effect of a shock. This property provides a bridge between millennial-scale dynamics and long-term climate transitions, ensuring that the projected trajectories remain robust across diverse time scales.
Data
Sources of Paleoclimate and Orbital Data
Figure 1: Evolution of climatic and orbital variables over the past 800,000 years. The panels display (from top to bottom) solar insolation at $60^{\circ}$N in June and in December, Antarctic temperature anomalies ($\Delta T$), and the logarithm of atmospheric CO2 concentration ($\ln C$), highlighting the strong phase-locking of climate variables to orbital cycles.
The empirical foundation of this study rests on three primary data series spanning the late Quaternary: atmospheric CO2 concentrations, Antarctic surface temperature anomalies, and solar insolation. These datasets provide the necessary temporal coverage to examine the long-term cointegrating relationships and the millennial-scale structural dynamics of the Earth system.
Atmospheric CO2 data are obtained from the composite ice-core record provided by the World Data Service for Paleoclimatology (Lüthi et al., 2008; Bereiter et al., 2015). This high-resolution dataset integrates measurements from the EPICA Dome C, Vostok, and other Antarctic ice cores, extending back to 800,000 years before present (yr BP). This record captures the natural variability of greenhouse gases across eight complete glacial-interglacial cycles, providing a robust baseline for evaluating climate sensitivity.
For temperature estimates, we utilize the Antarctic temperature anomaly record derived from stable isotope (deuterium) ratios in the EPICA Dome C ice core (Jouzel et al., 2007). The deuterium record provides a consistent proxy for regional Antarctic temperature, which serves as a reliable indicator for global climatic shifts due to the strong coupling between polar and global temperature variations.
Solar insolation data, representing the external orbital forcing, are computed using the IMCCE online numerical integration service based on the solution by Laskar et al. (2004). Specifically, we utilize the mean daily insolation at $60^{\circ}$N during the summer solstice in June. This metric is a standard choice (adopted by Lemoine, 2010) for capturing the orbital drivers of glacial-interglacial transitions, commonly referred to as Milankovitch forcing. The raw series of these variables, which exhibit clear cyclicality and phase-locking, are plotted in Figure 1.
Data Gridding and the Selection of Termination II
The raw ice-core samples are characterized by uneven temporal spacing, which typically becomes sparser with increasing age. To conduct a consistent time-series analysis using the Vector Error Correction Model (VECM), the data must be transformed into a regular temporal grid. We adopt a two-stage gridding process designed to preserve the integrity of the original physical signal while ensuring statistical reliability.
In the first stage, the raw data are allocated into 100-year bins. When a bin contains multiple observations, the values are averaged to reduce sampling noise and provide a representative value for that centennial interval. In the second stage, these 100-year records are aggregated into a 1,000-year grid, which serves as our primary unit of observation ($t$). For each 1,000-year grid point, the representative value is calculated by averaging the available 100-year bin values within that specific millennium.
Crucially, for our primary structural analysis, we do not employ interpolation during this aggregation process. If a 1,000-year window contains no underlying 100-year data points, it is intentionally left as a gap. This approach ensures that the resulting 1,000-year series remains a direct reflection of the observed ice-core record, free from any statistical artifacts that could be introduced by infilling missing values. Linear interpolation at the 100-year level is utilized exclusively for the Whole Period analysis (spanning 800,000 years) to facilitate a continuous estimation over the entire record where data sparsity is more pronounced.
As illustrated in Figure 2, our grid analysis identifies the presence of observational gaps across the late Quaternary. The longest continuous span that is entirely free of such gaps at the 1,000-year resolution—and therefore requires no interpolation whatsoever—perfectly coincides with the period known as Termination II (approximately 141,000 to 121,000 yr BP). This period captures the rapid transition from the penultimate glacial maximum to the Eemian interglacial.
By focusing our primary structural analysis on this gap-free Termination II sample, we ensure that our estimates of the Instantaneous Carbon Feedback (ICF) and the structural lag effects are derived from pure physical signals. This rigorous selection process maximizes the internal validity of our causal identification strategy, as it eliminates the possibility that the identified null Instantaneous Greenhouse Effect (IGE) is a product of data smoothing or interpolation bias.
Figure 2: Visualization of the data gridding structure and observational gaps. The upper and middle panels show the density of 100-year and 1,000-year grid points for Temperature and CO2, respectively. The bottom panel identifies data gaps (missing observations). The period corresponding to Termination II is highlighted as the longest continuous interval free of observational gaps for both variables.
Analysis
VECM Estimation Results
Table 1: Unit Root and Cointegration Test Results
Termination II (130-130k) Whole Period (800-800k)
ADF Unit Root Test $Z(t)$ $p$-value $Z(t)$ $p$-value
Temperature ($T$) 1.227-1.227 $0.662$ 3.303-3.303 $0.015$
[…]
Engle-Granger Test $Z(t)$ 10% Crit. $Z(t)$ 1% Crit.
Cointegration ($T$ on $\ln C$) 3.176-3.176* 3.077-3.077 7.871-7.871*** 3.910-3.910
Notes: ADF tests are performed with a constant. Critical values for the Engle-Granger test are based on MacKinnon (2010). ∗∗∗ and denote statistical significance at the 1% and 10% levels, respectively.
Table 2: Lag Order Selection Criteria
Termination II Whole Period
Lag FPE AIC HQIC SBIC FPE AIC HQIC SBIC
0 $0.0219$ $1.855$ $1.874$ $1.900$ $0.0222$ $1.868$ $1.873$ $1.880$
1 $0.0002$ 2.829-2.829 2.775-2.775 2.695-2.695 $0.0002$ 2.772-2.772 2.759-2.759 2.737-2.737
2 $0.0002$ 3.047-3.047 2.957-2.957 2.824-2.824 $0.0001$ 3.301-3.301 3.278-3.278 3.242-3.242
3 0.00010.0001* 3.157-3.157* 3.030-3.030* 2.845-2.845* $0.0001$ 3.375-3.375 3.343-3.343* 3.292-3.292*
4 $0.0002$ 3.108-3.108 2.945-2.945 2.707-2.707 0.00010.0001* 3.375-3.375* 3.334-3.334 3.269-3.269
Notes: * indicates lag order selected by the criterion. AIC: Akaike info criterion (FPE: final prediction error); SBIC: Schwarz Bayesian info criterion; HQIC: Hannan-Quinn info criterion.
Table 3: Comparison of VECM Results: Termination II vs. Whole Period
Termination II Whole Period
Panel A: Cointegrating Equation ($\beta$)
Variable Coefficient (s.e.) Coefficient (s.e.)
$T_{t-1}$ $1.000$ $1.000$
$\ln C_{t-1}$ 18.748-18.748*** $(2.312)$ 20.251-20.251*** $(1.081)$
$M_{t-1}$ $0.009$ $(0.038)$ 0.105-0.105*** $(0.017)$
Constant $103.300$ $164.974$
Panel B: Short-run Dynamics (Error Correction Model)
Independent Variables $\Delta T_{t}$ $\Delta\ln C_{t}$ $\Delta T_{t}$ $\Delta\ln C_{t}$
Error Correction ($\alpha$) 0.103-0.103** $0.002$ 0.164-0.164*** 0.0002-0.0002
$(0.042)$ $(0.001)$ $(0.017)$ $(0.001)$
$\Delta T_{t-1}$ $0.030$ $0.003$ 0.3460.346*** 0.0030.003***
$(0.096)$ $(0.003)$ $(0.034)$ $(0.001)$
$\Delta T_{t-2}$ 0.281-0.281*** $0.0004$ 0.123-0.123*** 0.0030.003**
$(0.093)$ $(0.003)$ $(0.033)$ $(0.001)$
$\Delta\ln C_{t-1}$ 15.20115.201*** 0.4470.447*** 6.0256.025*** 0.6000.600***
$(2.845)$ $(0.102)$ $(1.033)$ $(0.035)$
$\Delta\ln C_{t-2}$ 0.514-0.514 0.307-0.307*** $0.723$ 0.312-0.312***
$(3.008)$ $(0.108)$ $(1.080)$ $(0.037)$
$\Delta M_{t-1}$ 0.093-0.093*** 0.0005-0.0005 0.1460.146*** $0.0000$
$(0.033)$ $(0.001)$ $(0.020)$ $(0.001)$
Observations 129 799
AIC -3.498 -3.599
Notes: Standard errors are in parentheses. $T$, $\ln C$, and $M$ denote temperature, log-CO2, and orbital forcing, respectively. ∗∗∗, ∗∗, and denote significance at the 1%, 5%, and 10% levels.
Prior to estimating the structural dynamics of the climate system, we conduct a series of diagnostic tests to ensure the statistical validity of the Vector Error Correction framework. Table 1 reports the results of the Augmented Dickey-Fuller (ADF) unit root tests and the Engle-Granger cointegration tests. The ADF statistics indicate that both temperature and the logarithm of CO2 are non-stationary $I(1)$ processes across both the Whole Period and the Termination II samples. The Engle-Granger tests confirm the existence of a stable cointegrating relationship, providing the empirical basis for the VECM specification defined in Equation (1). Furthermore, the lag order selection criteria presented in Table 2 suggest that a second-order lag structure (where the level-VAR lag $k=3$ corresponds to VECM lag $p=2$) is optimal for capturing the millennial-scale dynamics while maintaining statistical parsimony.
The primary estimation results for the VECM are summarized in Table 3. Panel A presents the identified cointegrating vector ($\beta$), which defines the long-run physical bond between the variables. In both samples, we find a remarkably consistent and statistically significant equilibrium relationship. Specifically, for the Termination II sample—which provides the most reliable signal due to the absence of interpolation gaps—the relationship is estimated as $T-18.748\ln C=\textit{const}$. From this result, we derive the Earth System Sensitivity (ESS), defined as the total temperature response to a doubling of atmospheric CO2 concentration after all long-term feedbacks have reached equilibrium:
$\text{ESS}=18.748\times\ln(2)\approx 13.0\text{ K}$ (7)
Our estimate of 13.0 K is situated at the higher end of the range reported in the literature, such as the $9\pm 4$ K (95% CI: 7–13 K) estimate by Snyder (2016).
This high sensitivity can be understood through the conceptual framework of the PALAEOSENS project (Rohling et al., 2012). While the conventional Equilibrium Climate Sensitivity (ECS) typically accounts for “fast” feedbacks such as water vapor and clouds (often referred to as Charney sensitivity), our VECM, calibrated on the millennial-scale transitions of the Pleistocene, inherently captures the “slow” feedbacks of the Earth system. These include ice-sheet albedo changes, vegetation shifts, and long-term carbon-cycle responses that operate over multi-millennial time scales. According to Rohling et al. (2012), the inclusion of these slow feedbacks—which characterizes the transition from $S_{act}$ to $S_{ESS}$—can significantly amplify the total temperature response, potentially doubling the sensitivity compared to the Charney estimate. Our result of 13.0 K represents this comprehensive Earth System Sensitivity ($S_{ESS}$), reflecting the state-dependent amplification characteristic of major deglaciation events like Termination II.
Panel B of Table 3 delineates the short-run dynamics and the adjustment mechanisms of the system. The reduced-form lag coefficient for CO2 in the temperature equation, $(\Gamma_{1})_{12}=15.201$ ($p<0.01$), identifies the impact of CO2 changes from the previous millennium. Crucially, under the assumption of a null Instantaneous Greenhouse Effect ($a_{12}=0$)—a condition to be empirically validated in the subsequent section via the ICF Window for null IGE—this reduced-form parameter directly identifies the structural lag coefficient $(\Gamma_{1}^{*})_{12}$, representing the momentum of the greenhouse effect.
By recalling the structural identity $(I-A_{0})\Delta y_{t}=(I-A_{0})\alpha\beta^{\prime}y_{t-1}+\Gamma^{*}_{1}\Delta y_{t-1}+\dots+u_{t}$, the relationship between the structural and reduced-form coefficients is defined by the identity $\Gamma^{*}_{1}=(I-A_{0})\Gamma_{1}$. The specific element for the CO2-to-temperature lag can be expanded as follows:
$(\Gamma^{*}_{1})_{12}=(\Gamma_{1})_{12}-a_{12}(\Gamma_{1})_{22}-a_{13}(\Gamma_{1})_{32}$ (8)
Owing to the strict exogeneity of orbital parameters, past climate states cannot influence current orbital forcing, which dictates that the reduced-form coefficient $(\Gamma_{1})_{32}$ is effectively zero. Under this physical constraint, and imposing the null-IGE identification ($a_{12}=0$), the second and third terms in the expansion vanish. Consequently, the structural lag effect is directly identified by the estimated reduced-form coefficient: $(\Gamma^{*}_{1})_{12}=(\Gamma_{1})_{12}=15.201$.
This structural identification allows for a precise decomposition of the total temperature response observed in the 1,000-year window following a CO2 doubling event ($\Delta\ln C=\ln 2\approx 0.693$). The resulting warming is the sum of two distinct physical forces:
  1. Structural Momentum: The structural lag effect $(\Gamma^{*}_{1})_{12}$ provides an immediate "kick" of $15.201\times\ln(2)\approx 10.5$ K. This represents the powerful, delayed radiative feedback that manifests after the initial 1,000-year period of relative stability defined by the null-IGE identification.
  2. Equilibrium Adjustment: The error correction coefficient $\alpha=-0.103$ ($p=0.013$) quantifies the system’s restoring force toward the long-run ESS. Starting from the disequilibrium created by the CO2 innovation, this mechanism contributes an additional adjustment of approximately $13.0\text{ K}\times 0.103\approx 1.3$ K.
Combined, these forces generate a total warming of approximately 11.8 K within the first millennium after the lag period. The fact that this combination of delayed structural momentum and equilibrium pull so effectively accounts for the millennial-scale temperature variance explains why the model identifies $a_{12}\approx 0$ at the contemporaneous scale. In our framework, the greenhouse effect is not diminished; rather, it is structurally localized as a delayed feedback, allowing the contemporaneous driver of climatic shifts to be identified as orbital-induced temperature changes.
4.1.1 Consistency with Modern Anthropogenic Warming
The robustness of our SVECM estimates is further validated by comparing the model’s projected warming rate with observed modern temperature trends. Our results for Termination II indicate that a doubling of CO2 generates a total millennial response of approximately 11.8 K, comprising the 10.5 K structural momentum and the 1.3 K equilibrium adjustment. This translates to a centennial warming rate of 1.18 K per 100 years for a CO2 doubling.
To contextualize this within the modern era, we consider the increase in atmospheric CO2 since the pre-industrial period, from a baseline of approximately 280 ppm to the current level of roughly 428 ppm. In logarithmic terms, this corresponds to an accumulation of approximately 61% of a doubling:
$\text{Modern Accumulation Factor}=\frac{\ln(428/280)}{\ln(2)}\approx 0.61$ (9)
Applying our derived centennial rate to this factor, the model predicts a CO2-induced warming of:
$1.18\text{ K/century}\times 0.61\approx 0.72\text{ K/century}$ (10)
Remarkably, this paleo-derived estimate is in near-perfect alignment with the observed instrumental warming rate of approximately 0.73–0.80 K per century (equivalent to the observed 1.1–1.2 K rise over the last 150 years).
This high degree of correspondence suggests that the null-IGE identification ($a_{12}=0$) at the millennial scale does not diminish the long-term radiative importance of CO2. Instead, it reveals that the intrinsic sensitivity of the Earth system—as captured during the rapid transitions of the Pleistocene—is sufficient to account for the bulk of modern warming through delayed structural feedbacks and equilibrium adjustments. The fact that a model calibrated on 130,000-year-old ice-core data can so accurately recover the scale of 21st-century climate change provides powerful empirical support for our causal identification strategy.
Causality Allocation Analysis
4.2.1 Instantaneous Carbon Feedback (ICF)
The structural parameter $a_{21}$ is defined as a key element of the contemporaneous interaction matrix $A_{0}$ in our SVAR system, representing the simultaneous feedback within a single 1,000-year time step. Specifically, $a_{21}$ quantifies the Instantaneous Carbon Feedback (ICF), defined as the response of the logarithm of CO2 concentration to a change in temperature:
$a_{21}=\frac{\partial\ln C_{t}}{\partial T_{t}}$ (11)
To estimate this parameter from the paleoclimate record, we utilize a regression model formulated in first differences:
$\Delta\ln C_{t}=a_{21}\Delta T_{t}+v_{t}$ (12)
Since temperature and CO2 are characterized as $I(1)$ processes, their first-differenced series are stationary, or $I(0)$. This transformation ensures that the estimation is not susceptible to the problem of spurious regression, which often plagues level-based analyses of persistent time series.
However, estimating Equation (12) via Ordinary Least Squares (OLS) is potentially problematic due to the inherent endogeneity of the climate-carbon system. Given the simultaneous nature of the coupling, a reverse causal link also exists:
$\Delta T_{t}=a_{12}\Delta\ln C_{t}+w_{t}$ (13)
where $a_{12}$ represents the Instantaneous Greenhouse Effect (IGE). If a significant IGE exists at the millennial resolution, the OLS estimate for $a_{21}$ would suffer from simultaneity bias, as $\Delta T_{t}$ would be correlated with the error term $v_{t}$.
To address this concern and ensure the consistency of our estimates, we first employ Instrumental Variables (IV) estimation. Our identification strategy utilizes Milankovitch orbital parameters, specifically the winter solstice insolation at 60N ($M^{\text{N60D}}_{t}$), and lagged temperature levels ($T_{t-2}$) as instruments. These celestial mechanics provide a source of exogenous variation in temperature that is physically independent of the Earth’s internal carbon cycle. By utilizing these instruments, we isolate the variations in temperature that are driven by external forcing, thereby identifying the pure physical response of atmospheric CO2 to climatic warming.
The estimation results are summarized in Table 4. Crucially, the endogeneity test (a robust Durbin-Wu-Hausman test) for the Termination II sample yields a $p$-value of 0.864. This result indicates that we cannot reject the null hypothesis that the temperature change is exogenous with respect to CO2 innovations. In the context of our causality allocation, this provides strong empirical evidence that the contemporaneous reverse causality from CO2 to temperature (the IGE) is statistically undetectable at the 1,000-year resolution. Consequently, OLS estimation is not only consistent but also more efficient than IV estimation. The resulting OLS estimate for Termination II is $\hat{a}_{21}=0.02082$ ($z=8.04,p<0.01$), which we adopt as the representative structural parameter for the ICF.
Table 4: IV Estimation Results for the Instantaneous Carbon Feedback ($a_{21}$)
Sample Period Estimate $z$ 95% Conf. Interval Sensitivity
$($$a_{21}$) Lower Upper (ppm/K)
Whole Period $0.0222$ $5.44$ $0.0142$ $0.0302$ $6.22$
Termination II $0.0219$ $3.81$ $0.0106$ $0.0332$ $6.13$
Termination II (OLS) $0.02082$ $8.04$ $0.0157$ $0.0259$ $5.82$
Diagnostics for IV Estimation 1st $F$ Overid. Endog. Obs.
Whole Period $28.48$ $0.963$ $0.005$ 801
Termination II $11.91$ $0.054$ $0.864$ 130
Notes: The dependent variable is $\Delta\ln C_{t}$ and the endogenous regressor is $\Delta T_{t}$ as defined in Equation (12). Instruments for the Whole Period are $\{M^{\text{N60J}}_{t},M^{\text{N60D}}_{t}\}$ and for Termination II are $\{M^{\text{N60D}}_{t},T_{t-2}\}$. Overid. reports the $p$-value for the Hansen J or Sargan tests. Endog. reports the $p$-value for the Durbin-Wu-Hausman endogeneity test; the high $p$-value for Termination II (0.864) indicates that the null hypothesis of exogeneity for $\Delta T_{t}$ cannot be rejected, justifying the use of the OLS estimate ($\hat{a}_{21}=0.02082$). The convergence between OLS and IV estimates supports the identification of a null instantaneous greenhouse effect ($a_{12}=0$). Sensitivity is converted as $a_{21}\times 280$ ppm/K for comparison.
4.2.2 Instantaneous Greenhouse Effect (IGE)
The absence of a significant contemporaneous IGE ($a_{12}=0$), as suggested by the endogeneity diagnostics, can be further explored and validated through the Causality Allocation Curve (CAC). This curve represents the set of ($a_{12},a_{21}$) pairs that are consistent with the estimated reduced-form covariance matrix of the VECM residuals, assuming the orthogonality of structural innovations.
As illustrated in Figure 3, the CAC defines the structural trade-off between the allocated strengths of IGE and ICF. A remarkable consistency is observed between two independent analytical paths. The ICF value identified by our first-difference OLS ($\hat{a}_{21}=0.02082$) almost perfectly coincides with the structural $a_{21}$ value where the CAC intersects the $a_{12}=0$ axis ($\hat{a}_{21}=0.02084$), which was estimated independently via the Method of Moments based on the system-wide VECM residuals (as reported in Table 5).
This convergence between instrumental variable diagnostics and system-wide residual analysis is a significant finding. It suggests that the model, including the VECM lag structure and the null-IGE identification ($a_{12}=0$), captures the underlying physical dynamics of the Pleistocene with high fidelity. If the model had been mis-specified, or if a significant instantaneous greenhouse effect were truly present, the OLS estimate would be expected to diverge significantly from the point where $a_{12}=0$.
To quantify the statistical stability of this identification, we define the ICF Window for null IGE as the range of ICF values for which the null hypothesis of $a_{12}=0$ cannot be rejected. As shown in Panel B of Table 5, a sensitivity analysis conducted at the boundaries of the 95% confidence interval for our ICF estimate confirms that the IGE remains statistically indistinguishable from zero across the entire plausible range of carbon cycle feedback.
Furthermore, we consider the median sensitivity estimate from Frank et al. (2010), which corresponds to $a_{21}\approx 0.0275$ when converted to our logarithmic scale (7.7 ppm/K divided by a 280 ppm baseline). While this value sits near the boundary of the ICF Window for null IGE (being rejected at the 5% level but not at the 1% level), our data-driven estimate of 0.0208 lies precisely at its center. This indicates that while the consensus in the carbon-cycle literature is broadly compatible with a temperature-led causal structure, our empirical identification achieves a state of maximal statistical consistency. These results collectively reinforce the conclusion that the greenhouse effect in the paleoclimate is not an instantaneous phenomenon but is instead manifested through the structural lagged effects and long-run equilibrium adjustments described in the previous section.
Table 5: $M$ Estimation Results for ICF Window
Restriction / Parameter Estimate $z$ $P>|z|$ 95% Conf. Interval
Lower Upper
Panel A: Estimating $a_{21}$ assuming null marginal greenhouse forcing ($a_{12}=0$)
$a_{21}$ (Carbon Feedback) $0.02084$ $7.81$ $0.000$ $0.0156$ $0.0261$
Panel B: Sensitivity analysis of $a_{12}$ at the boundaries of the $a_{21}$ 95% CI
$a_{12}$ (at $a_{21}=0.0156$) $5.3805$ $2.00$ $0.046$ $0.1052$ $10.6557$
$a_{12}$ (at $a_{21}=0.0261$) 6.9366-6.9366 1.71-1.71 $0.087$ 14.8738-14.8738 $1.0006$
Notes: Estimates are obtained using the Method of Moments (MM) based on VECM residuals. Panel A identifies the ICF value required to satisfy the condition of a null instantaneous greenhouse effect ($a_{12}=0$). Panel B demonstrates the statistical stability of the IGE estimate at the boundaries of the ICF’s 95% confidence interval. The resulting ICF Window for null IGE encompasses our empirical OLS estimate, confirming the structural consistency of the model.
Figure 3: The Causality Allocation Curve and ICF Window for null IGE. The curve illustrates the structural trade-off between the Instantaneous Greenhouse Effect (IGE, $a_{12}$) and the Instantaneous Carbon Feedback (ICF, $a_{21}$). The white region identifies the ICF Window for null IGE, where the hypothesis $a_{12}=0$ is not rejected at the 1% level. Our OLS estimate for Termination II ($\hat{a}_{21}=0.02082$) lies centrally within this window, highlighting the robustness of the null-IGE identification.
Impulse Response Simulation of Climate-Carbon Dynamics
In this final part of the analysis, we evaluate the dynamic stability and the long-term sensitivity of the identified Earth system via impulse response functions (IRFs). By utilizing the moving average representation of the estimated SVECM, we simulate the propagation of structural innovations across millennial time scales. These simulations provide a visual and quantitative synthesis of the Instantaneous Carbon Feedback (ICF), the Instantaneous Greenhouse Effect (IGE), and the structural lagged effects identified in the preceding sections.
Crucially, the IRF trajectories allow us to distinguish between the immediate structural impacts and the long-term adjustment toward the ESS discussed in Section 4.1. As we move from the instantaneous response (governed by $a_{ij}$) toward the multi-millennial equilibrium, the transition captures the progression from the initial radiative forcing to the fully realized Earth System Sensitivity. This dynamic visualization aligns with the perspective of Rohling et al. (2012), who emphasize that the measured sensitivity is a function of the time scale considered, with the full magnitude of slow feedbacks only becoming apparent as the system approaches its new steady state.
4.3.1 System Response to a CO2 Doubling Innovation
We first simulate a scenario representing an abrupt doubling of atmospheric CO2 concentration. This is implemented as a structural innovation in the carbon cycle, $u_{2t}=\ln(2)\approx 0.693$, applied to a system initially at a stable equilibrium (baseline CO2 = 280 ppm). The resulting trajectories for temperature and CO2 are presented in Figure 4.
As shown in the left panels of Figure 4, the model captures a distinct two-stage temperature response. During the first 1,000-year interval immediately following the shock, the temperature anomaly remains effectively zero. This period of stasis is a direct consequence of our null-IGE identification ($a_{12}=0$), which implies that CO2 concentrations do not exert an additional thermal impact within the same millennial time step.
However, as the system enters the second millennium, we observe a sharp and powerful rise in temperature. This "thermal breakout" is driven by the structural lag coefficient $(\Gamma^{*}_{1})_{12}=15.201$, which delivers a structural momentum of approximately 10.5 K. Simultaneously, the error correction mechanism ($\alpha=-0.103$) begins to pull the system toward its long-run Earth System Sensitivity (ESS). Over the subsequent 10,000 years, the temperature continues to rise as the slow-acting feedbacks of the Pleistocene climate reach maturity. As indicated by the long-run impact matrix $\Xi$, the temperature eventually stabilizes at a permanent increase of approximately 13.0 K, consistent with the ESS derived from the cointegrating vector.
This simulation illustrates that the greenhouse effect in the paleoclimate is not an instantaneous force but a structurally delayed phenomenon. The ability of the model to account for the observed variance in temperature through this delayed momentum, while maintaining $a_{12}=0$, provides a compelling physical explanation for the phase-locking observed in the ice-core records.
Figure 4: Impulse response functions for a CO2 doubling shock ($u_{2t}=\ln 2$). The plots show the evolution of CO2 concentration (upper panels) and Antarctic temperature anomalies (lower panels) over a 100,000-year horizon (left) and a detailed 10,000-year view (right). Note the 1,000-year lag in the temperature response, followed by a transition toward the long-run ESS of 13.0 K.
Figure 5: Impulse response functions for a 2 K temperature shock ($u_{1t}=2.0$). The plots illustrate the immediate response of the carbon cycle via the ICF and the long-term permanent shift in both variables. The secondary rise in temperature following the initial shock demonstrates the operation of the identified positive feedback loop between the thermal and carbon cycles.
4.3.2 System Response to a 2 K Temperature Shock
To evaluate the efficiency of the ICF and the robustness of the carbon cycle, we simulate a structural innovation in temperature of $u_{1t}=2.0$ K. This scenario, depicted in Figure 5, represents an abrupt warming event independent of the greenhouse effect, such as a shift in orbital forcing or ocean circulation.
Upon the occurrence of the shock, the CO2 concentration immediately begins to rise, driven by the ICF coefficient $a_{21}=0.02082$. At the pre-industrial baseline of 280 ppm, this corresponds to an instantaneous release of approximately 5.8 ppm per degree of warming. In the initial 1,000-year step, the 2 K shock induces a rise of approximately 11.7 ppm in atmospheric CO2.
The long-term trajectory shown in the left panels of Figure 5 reveals the operation of the cointegrating bond. As the temperature anomaly is sustained, the carbon cycle continues to adjust toward its new equilibrium level. Due to the property of cointegration in a non-stationary system, a temporary structural shock to temperature leads to a permanent shift in the CO2 baseline. The CO2 level eventually reaches a new steady state, representing the integrated response of ocean degassing and terrestrial biosphere adjustments.
Crucially, the temperature itself exhibits a slight secondary increase after the initial shock. This is the structural representation of the positive feedback loop: the initial warming triggers a release of CO2 (ICF), which in turn amplifies the warming through the structural lag effect in subsequent millennia. The stabilization of the system at a higher baseline for both variables confirms that the identified VECM is dynamically stable and that the cointegrating relationship $T-18.748\ln C=\text{const}$ acts as a robust anchor for the Earth system’s long-term behavior.
Concluding Remarks
In this study, we have developed and estimated a Structural Vector Error Correction Model (SVECM) to disentangle the complex, bidirectional causalities between temperature and atmospheric CO2 concentrations during the Pleistocene. By focusing on the high-quality, gap-free record of Termination II and utilizing the strict exogeneity of orbital forcing for structural identification, we have provided a new perspective on the Earth system’s sensitivity and its response to carbon innovations.
The primary finding of our analysis is the identification of a null Instantaneous Greenhouse Effect (IGE) at the millennial resolution. Our results consistently indicate that the contemporaneous parameter $a_{12}$ is statistically indistinguishable from zero. This does not imply that the greenhouse effect is physically absent; rather, it reveals that the radiative impact of CO2 is structurally localized as a delayed feedback. Our model identifies a powerful structural lag effect, $(\Gamma^{*}_{1})_{12}$, which delivers a warming of approximately 10.5 K in the millennium following a CO2 doubling. When combined with an equilibrium adjustment of approximately 1.3 K toward a long-run Earth System Sensitivity (ESS) of 13.0 K, the total millennial response reaches approximately 11.8 K.
Crucially, the identification of $a_{12}=0$ is not a mere assumption but an empirical necessity derived from the data. The Instantaneous Carbon Feedback (ICF), estimated via first-difference OLS as $\hat{a}_{21}=0.02082$, lies at the precise center of the ICF Window for null IGE. The convergence between independent analytical paths—instrumental variable diagnostics for the carbon cycle and system-wide residual analysis for the thermal cycle—provides definitive empirical verification for the validity of our structural identification. While the median sensitivity from previous ensemble reconstructions such as Frank et al. (2010) sits at the boundary of statistical rejection for a null IGE, our data-driven estimate achieves maximal statistical consistency with the observed millennial dynamics.
Furthermore, the robustness of our paleo-derived parameters is demonstrated by their remarkable predictive power for the modern era. The projected warming rate of 0.72 K per century, derived solely from our Termination II SVECM and adjusted for the modern CO2 accumulation factor of 0.61, is in near-perfect alignment with the instrumental observation of 0.73–0.80 K per century. The fact that a model calibrated on 130,000-year-old ice-core data can so accurately recover the scale of 21st-century climate change suggests that the structural properties of the Earth system’s coupling have remained fundamentally consistent across diverse climatic regimes.
Our impulse response simulations further clarify the physical mechanism: the Earth system operates as a coupled engine where temperature changes, driven by orbital or other external shocks, immediately trigger a carbon response through the ICF. This released CO2 then acts as a powerful long-term regulator, driving the planet toward a new climatic baseline through delayed structural momentum and long-run equilibrium adjustment.
In conclusion, this research bridges the gap between paleoclimatic structural analysis and modern climate monitoring. By identifying the specific temporal localization of the greenhouse effect, we provide a framework that explains both the millennial-scale phase-locking observed in the ice cores and the centennial-scale warming observed in the instrumental record. These findings underscore the critical importance of accounting for delayed feedbacks in long-term climate projections and demonstrate the enduring value of the Pleistocene record in deciphering the future of our planet.
Appendix: Theoretical Background of VECM and Impulse Response
Decay of Shocks in a Stationary VAR
Consider the following VAR($p$) model:
$y_{t}=\Phi_{1}y_{t-1}+\Phi_{2}y_{t-2}+\dots+\Phi_{p}y_{t-p}+\epsilon_{t}$ (14)
When the stability condition [^2] All roots of the characteristic equation $\text{det}(I-\Phi_{1}z-\dots-\Phi_{p}z^{p})=0$ lie outside the unit circle in the complex plane. In this case, $z=1$ is not a root, and $\text{det}(G=I-\sum\Phi_{i})\neq 0$ holds. Such a system is also referred to as being stationary. is satisfied, this equation can be transformed into the following MA($\infty$) representation:
$y_{t}=\Psi_{0}\epsilon_{t}+\Psi_{1}\epsilon_{t-1}+\Psi_{2}\epsilon_{t-2}+\dots=\sum_{h=0}^{\infty}\Psi_{h}\epsilon_{t-h}\quad(\Psi_{0}=I)$ (15)
Here, the coefficient matrix $\Psi_{h}$ represents the impact (impulse response) of a current shock $\epsilon_{t}$ on the level $y_{t+h}$ after $h$ periods, and it is defined recursively as $\Psi_{h}=\Phi_{1}\Psi_{h-1}+\dots+\Phi_{p}\Psi_{h-p}$.
Assuming a constant shock has been applied from the past to the present ($\epsilon_{t-h}=\epsilon$), Equation (15) yields $y_{t}=(\sum_{h=0}^{\infty}\Psi_{h})\epsilon$. On the other hand, assuming a long-run equilibrium state ($y_{t}=y_{t-1}=\dots=y$) and substituting this into Equation (14), we obtain $\left[I-\sum_{i=1}^{p}\Phi_{i}\right]y=Gy=\epsilon$. By comparing the two, the following relationship is derived:
$\sum_{h=0}^{\infty}\Psi_{h}=G^{-1}\quad(\text{Long-run Multiplier})$ (16)
Under the stability condition, the right-hand side (long-run multiplier) is determined as a finite value, which requires the series on the left-hand side to converge to zero. As a necessary condition, the impulse response must vanish over time ($\lim_{h\to\infty}\Psi_{h}=0$).
Redefinition via Differencing
To encompass non-stationary cases, we redefine the level response $\Psi_{h}$ as the cumulative sum of responses in differences. When the difference $\Delta y_{t}=y_{t}-y_{t-1}$ is stationary, $\Delta y_{t}$ has the following MA representation:
$\Delta y_{t}=C_{0}\epsilon_{t}+C_{1}\epsilon_{t-1}+C_{2}\epsilon_{t-2}+\dots=\sum_{j=0}^{\infty}C_{j}\epsilon_{t-j}\quad(C_{0}=I)$ (17)
where $C_{j}=\Psi_{j}-\Psi_{j-1}$ represents the impact of the shock $\epsilon_{t-j}$ on the current "change" $\Delta y_{t}$. Given that the level $y_{t}$ is the cumulative sum of past differences $y_{t}=y_{0}+\sum_{j=1}^{t}\Delta y_{j}$, the impulse response for the level, $\Psi_{h}=C_{h}+\Psi_{h-1}$, can be rewritten as:
$\Psi_{h}=\sum_{j=0}^{h}C_{j}$
Application to Non-stationary Cases
In systems where the characteristic equation has a unit root ($z=1$), the matrix $G$ becomes non-invertible ($\det(G)=0$), and the long-run multiplier in Equation (16) cannot be defined. Here, we treat the case where there is exactly one unit root, and all other roots lie outside the unit circle, as the non-stationary case. Such non-stationary series are referred to as $I(1)$, and their differenced series are stationary.
If the differenced series $\Delta y_{t}$ is stationary, then $\sum_{j=0}^{\infty}C_{j}$ in (17) becomes a finite value. Therefore, $C_{j}$ must converge to zero ($\lim_{j\to\infty}C_{j}=0$). Furthermore, in this case, $\Psi_{h}$ converges to a non-zero constant matrix $\Xi$:
$\Psi_{\infty}=\lim_{h\to\infty}\Psi_{h}=\sum_{j=0}^{\infty}C_{j}=\Xi\neq 0$ (18)
This $\Xi$ refers to the permanent effect (Permanent Effect) of a temporary shock on the absolute level of the variables. In summary, while the impact of a shock in a stationary VAR is transitory and returns to zero, in a non-stationary system such as VECM, the impact on the "changes" accumulates and persists, leading to a permanent shift in the baseline level of the system.
The impulse response after $h$ periods can be decomposed as:
$\Psi_{h}=\sum_{j=0}^{h}C_{j}=\left(\sum_{j=0}^{\infty}C_{j}\right)-\left(\sum_{j=h+1}^{\infty}C_{j}\right)=\Xi+\Upsilon_{h}$
From Equation (18), it is evident that $\lim_{h\to\infty}\Upsilon_{h}=\lim_{h\to\infty}(\Psi_{h}-\Xi)=0$, meaning $\Upsilon_{h}$ represents the decaying transitory component.
Case with Cointegration: The Resilience of the Physical Bond
The VAR($p$) model in Equation (14) can be reformulated as the following VEC model:
$\Delta y_{t}=\Pi y_{t-1}+\Gamma_{1}\Delta y_{t-1}+\dots+\Gamma_{p-1}\Delta y_{t-p+1}+\epsilon_{t}$ (19)
where $\Pi=\alpha\beta^{\prime}$ represents the long-run equilibrium (cointegrating) relationships. Note that $\Phi_{1}=I+\Pi+\Gamma_{1}$, $\Phi_{i}=\Gamma_{i}-\Gamma_{i-1}$, and $\Phi_{p}=-\Gamma_{p-1}$. Engle and Granger (1987) demonstrated that for $I(1)$ series, a linear combination $\beta^{\prime}y_{t}$ exists that is stationary ($I(0)$).
A critical implication of cointegration is the orthogonality between the permanent effect $\Xi$ and the cointegrating vector $\beta$:
$\beta^{\prime}\Xi=0$ (20)
This mathematical property carries a profound physical meaning: while an exogenous shock (such as a sudden change in orbital forcing or a carbon cycle innovation) can permanently shift the absolute levels of temperature and CO2, it is fundamentally incapable of breaking the equilibrium "bond" between them. The variables move to a new climatic baseline, but the underlying physical relationship, defined by $\beta$, remains intact.
Johansen (1991) provided the explicit representation for this permanent impact matrix:
$\Xi=\beta_{\perp}(\alpha_{\perp}^{\prime}\Gamma\beta_{\perp})^{-1}\alpha_{\perp}^{\prime}$ (21)
where $\Gamma=I-\sum_{i=1}^{p-1}\Gamma_{i}$. This expression confirms that the permanent shift $\Xi$ occurs only within the space spanned by $\beta_{\perp}$ (the orthogonal complement of $\beta$), ensuring that the long-run impact never deviates from the cointegrating attractor. Thus, the condition $\mathbb{E}[\beta^{\prime}y_{t}]=0$ is maintained even as the Earth system undergoes permanent transitions between glacial and interglacial states.

### v3

@@
-Structural Cointegration of the Paleoclimate
+arXiv:2601.06986v3[physics.ao-ph] 28 Feb 2026
 Abstract
-This paper investigates the coupled dynamics of temperature and atmospheric CO2 concentrations during Pleistocene glacial cycles using a Structural Vector Error Correction Model (SVECM).
-To address the controversy regarding causal leads and lags, we develop a structural identification framework that utilizes Milankovitch orbital parameters as instrumental variables to disentangle the Instantaneous Greenhouse Effect (IGE) and the Instantaneous Carbon Feedback (ICF) at a millennial resolution.
-Focusing on the gap-free interval of Termination II, we employ a Causality Allocation Curve (CAC) derived from system-wide residuals to identify the causal structure that maximizes statistical consistency with the observed record.
-This approach allows for the estimation of Earth System Sensitivity (ESS) based purely on observational data and the decomposition of warming dynamics into structural momentum and equilibrium adjustment.
-By demonstrating the consistency of these paleo-derived structural parameters with modern instrumental records, this study establishes a robust econometric foundation for integrating long-term climate analysis with contemporary climate monitoring.
+Using a gap-free, millennial-resolution record of atmospheric CO2 and Antarctic temperature since the Last Interglacial (130 ka ~), we structurally identify the feedback architecture of the climate–carbon system.
+The series are cointegrated, supporting estimation with a Vector Error Correction Model (VECM).
+The cointegrating relationship implies a temperature response of 13.0 K per CO2 doubling.
+Structural identification combining a Causality Allocation Curve (CAC) and Milankovitch-cycle instrumental variables yields a contemporaneous carbon response (CCR) of 5.77 ppm/K, while the contemporaneous temperature response (CTR) is statistically indistinguishable from zero.
+Accounting for lagged feedbacks, the cumulative temperature response within one millennium of CO2 doubling reaches 11.8 K.
 Introduction
-The strong coupling between atmospheric CO2 concentrations and global temperatures over the glacial-interglacial cycles of the Quaternary period is one of the most striking features of the Earth’s climate system.
-High-resolution ice-core records from Antarctica, spanning the last 800,000 years, have provided unequivocal evidence that CO2 and temperature covary with remarkable precision across multiple terminations and glaciations (Jouzel et al., 2007; Lüthi et al., 2008).
-This tight correlation forms the empirical basis for our understanding of the greenhouse effect as a fundamental regulator of the Earth’s long-term climatic state.
-However, the precise causal architecture of this relationship remains a subject of intense debate in paleoclimatology.
-Early investigations into the timing of these transitions suggested that Antarctic temperature increases preceded CO2 rises by several centuries during glacial terminations, leading to the hypothesis that CO2 acts primarily as an amplifier of orbitally-induced warming rather than its primary driver (Caillon et al., 2003).
-Conversely, recent re-evaluations of ice-core gas-age chronologies have argued that the perceived lag may be a statistical artifact of the firn densification process, suggesting that CO2 and temperature are effectively synchronized at the millennial scale (Parrenin et al., 2013).
-This “lead-lag” controversy is not merely a chronological puzzle; it strikes at the heart of our ability to quantify the true climate sensitivity and the structural feedbacks inherent in the Earth system.
-The fundamental difficulty in resolving this causality lies in the simultaneous nature of the climate-carbon coupling.
-Within the observational resolution of the paleoclimatic record, temperature and CO2 operate in a continuous feedback loop: warming triggers the release of CO2 from the oceans and terrestrial biosphere, while increased CO2 further warms the planet through radiative forcing.
-In econometric terms, this constitutes a problem of simultaneity and endogeneity, where standard correlation analyses or simple Granger-causality tests are insufficient to isolate the structural parameters of the system.
-Without a robust identification strategy, it is impossible to determine whether the observed covariance is driven by an instantaneous greenhouse effect or by a temperature-driven carbon response.
-In this paper, we address this identification challenge by applying a Structural Vector Error Correction Model (SVECM) to the Pleistocene ice-core record.
-Our approach builds upon the methodology of structural macroeconometrics, utilizing the strict exogeneity of Earth’s orbital parameters—Milankovitch forcing—to achieve causal identification.
-We distinguish between two contemporaneous mechanisms operating within a 1,000-year time step: the Instantaneous Greenhouse Effect (IGE), representing the immediate thermal response to CO2 innovations, and the Instantaneous Carbon Feedback (ICF), representing the carbon cycle’s response to warming.
-By analyzing the gap-free record of Termination II, we demonstrate that the Pleistocene climate is characterized by a null Instantaneous Greenhouse Effect ($a_{12}\approx 0$) at the millennial resolution.
-Instead, we show that the greenhouse effect is structurally localized as a powerful delayed feedback.
-Our model identifies a structural lag of 1,000 years, after which a CO2 doubling triggers a cumulative temperature increase of approximately 11.8 K, comprising a 10.5 K structural momentum and a 1.3 K long-run equilibrium adjustment.
-Crucially, we validate these paleo-derived structural parameters by projecting them onto the modern era.
-We show that the centennial warming rate derived from our Pleistocene model accurately recovers the observed instrumental warming of the last 150 years.
-This consistency suggests that the fundamental coupling between the thermal and carbon cycles has remained stable across vastly different climatic regimes.
-By providing a definitive empirical verification of the structural lag in the greenhouse effect, this study offers a new framework for reconciling the millennial-scale dynamics observed in the ice cores with the centennial-scale warming observed in the modern anthropogenic era.
-While earlier studies, such as Kaufmann et al.
-(2006), have successfully applied VECM to instrumental climate records to identify long-run relationships, they relied on pre-defined radiative forcing formulas and reduced-form estimations.
-By contrast, our study extends this econometric framework into the structural domain (SVECM).
-By utilizing the orbital forcing as a source of exogenous variation, we identify the distinct structural parameters of the Earth system—IGE, ICF, and the structural lag effects.
-This allows us to estimate an Earth System Sensitivity (ESS) that is derived purely from observational data, consistent with the long-term feedback framework proposed by the PALAEOSENS project (Rohling et al., 2012).
+The tight coupling between Antarctic temperature and atmospheric carbon dioxide (CO2) concentrations over the late Quaternary is one of the most striking features of the Earth system.
+While the high degree of correlation between these variables is well-established, the nature of their contemporaneous interaction—specifically the direction, magnitude, and timing of feedbacks at millennial resolutions—remains a subject of intense debate.
+Establishing a robust causal architecture is not only vital for understanding past glacial-interglacial transitions but also for constraining long-term Earth System Sensitivity (ESS) in the context of anthropogenic warming.
+The statistical analysis of paleoclimate records faces three fundamental challenges.
+First, the high persistence of temperature and CO2 series carries the risk of spurious regression (Kaufmann et al., 2006; Cummins et al., 2022).
+Because these variables exhibit the properties of non-stationary, $I(1)$ processes, standard regression techniques can yield misleadingly significant correlations that lack physical causality.
+Second, a methodological schism exists regarding the appropriate framework for these series; while some advocate for a cointegration approach (Kaufmann et al., 2006; Miller, 2019), others argue that the system is ultimately mean-reverting and stationary, favoring standard Vector Autoregression (VAR) models (Davidson et al., 2016).
+Third, the irregular temporal spacing of ice-core proxies often necessitates linear interpolation, which has been shown to introduce substantial size distortions and biases in statistical testing (Miller, 2019).
+Beyond these statistical hurdles, a long-standing climate-physical debate persists regarding the “lead-lag” relationship between temperature and CO2.
+Early studies suggested that CO2 lagged Antarctic temperature by several centuries during deglaciations (Caillon et al., 2003), implying that carbon was a follower rather than a driver of initial warming.
+More recent re-evaluations, accounting for gas-ice age uncertainties, have proposed a nearly synchronous relationship (Parrenin et al., 2013).
+However, these findings often rely on reduced-form correlations that cannot isolate structural innovations from reciprocal feedbacks.
+While Lemoine (2010) utilized orbital parameters as instrumental variables (IV) to identify a significantly positive—yet relatively modest—climate–carbon feedback factor, the short-run structural dynamics within a unified millennial-scale framework remain under-explored.
+This study addresses these challenges by applying a Structural Vector Error Correction Model (SVECM) to a high-resolution, gap-free ice-core dataset.
+Our contribution is twofold.
+First, we identify a specific 130,000-year interval (the 130 ka sample) that is entirely free of observational gaps at a 1,000-year resolution.
+By utilizing this interval, we eliminate the need for interpolation, thereby physically neutralizing the statistical artifacts warned of by Miller (2019).
+Second, we employ a structural identification strategy that utilizes the strict exogeneity of Milankovitch cycles as an instrument to disentangle the contemporaneous temperature response (CTR, $a_{12}$) from the contemporaneous carbon response (CCR, $a_{21}$).
+Our results reveal a stark asymmetry in the contemporaneous climate–carbon interaction.
+We find a robust, statistically significant CCR ($\hat{a}_{21}>0$), consistent with the magnitude suggested by last-millennium reconstructions (Frank et al., 2010).
+In contrast, the CTR ($a_{12}$) is found to be statistically indistinguishable from zero at the millennial scale.
+This identification implies that while temperature changes trigger an immediate carbon release, the radiative feedback of CO2 onto temperature requires structural lags to manifest.
+This framework provides a unified explanation for the perceived “synchronicity” in deglacial records as a manifestation of the CCR, while reconciling the “lag” observed in other contexts as the physical adjustment time of the thermal system.
+By bridging the gap between time-series econometrics and paleoclimatic physics, we provide a more precise characterization of the feedbacks that govern the Earth’s long-term climate sensitivity.
 Model
@@
 To circumvent this risk while preserving the critical information contained in the levels of the variables, we employ a Vector Error Correction Model (VECM).
-This framework is appropriate when variables are cointegrated, meaning that while individual series may drift non-stationarily, a specific linear combination of them remains stationary.
-This stationary combination represents a long-run equilibrium or a physical bond that has persisted through numerous glacial-interglacial cycles.
-The VECM is formulated as follows:
+This framework is appropriate when variables are cointegrated, meaning that while individual series may drift non-stationarily, a stable linear combination exists.
+In this study, we consider a three-variable system $y_{t}=(T_{t},\ln C_{t},M_{t})^{\prime}$, where $T_{t}$ is Antarctic temperature, $\ln C_{t}$ is the logarithm of CO2 concentration, and $M_{t}$ represents external orbital (Milankovitch) forcing.
+The VECM is expressed as:
 $\Delta y_{t}=\alpha\beta^{\prime}y_{t-1}+\Gamma_{1}\Delta y_{t-1}+\dots+\Gamma_{p-1}\Delta y_{t-p+1}+\epsilon_{t}$ (1)
-where $y_{t}=(T_{t},\ln C_{t},M_{t})^{\prime}$ is the vector of endogenous variables representing temperature ($T_{t}$), log-transformed CO2 concentration ($\ln C_{t}$), and orbital forcing ($M_{t}$).
-$\Delta$ denotes the first-difference operator, and $\epsilon_{t}$ is a vector of reduced-form residuals.
-The matrix $\Pi=\alpha\beta^{\prime}$ contains the essential information regarding the cointegrating space.
-The vector $\beta$ defines the long-run equilibrium toward which the climate system gravitates over centennial to millennial scales.
-The existence of a stable $\beta$ suggests a fundamental consistency in the coupling of the thermal and carbon cycles across diverse climatic regimes.
-The matrix $\alpha$ represents the speed of adjustment or the restoring force of the system.
-When a stochastic shock—such as an orbital perturbation—displaces the system from its equilibrium path, the error correction term ($\beta^{\prime}y_{t-1}$) becomes non-zero.
-The coefficient $\alpha$ dictates how rapidly the system initiates corrective dynamics.
-This mechanism constitutes the mathematical representation of the climate system’s self-regulation.
+where $\beta$ contains the cointegrating vector, $\alpha$ represents the adjustment speeds to equilibrium, and $\Gamma_{i}$ captures the short-run lagged dynamics.
+The error term $\epsilon_{t}$ denotes the reduced-form innovations, which are generally correlated.
 Structural Identification and Causality Allocation
-To isolate structural causalities from the reduced-form correlations, we assume that the VECM is derived from an underlying Structural VAR (SVAR) model.
-This approach follows the framework established by Sims (1980), which provides a basis for analyzing endogeneity and structural innovations within a system of interrelated time series by utilizing the residuals of the estimated model.
-By applying this structural methodology to the paleoclimatic system, we can define a model that explicitly accounts for the contemporaneous interactions between temperature and carbon cycles:
-$(I-A_{0})\Delta y_{t}=(I-A_{0})\alpha\beta^{\prime}y_{t-1}+\Gamma^{*}1\Delta y{t-1}+\dots+u_{t},$ $u_{t}\sim\text{IID}(0,I)$ (2)
-The matrix $A_{0}$ represents the simultaneous feedbacks within the same 1,000-year time step. The reduced-form residuals $\epsilon_{t}$ are related to the structural shocks $u_{t}$ via the relation:
-$u_{t}=(I-A_{0})\epsilon_{t}$ (3)
-Here, the elements $a_{ij}$ of the structural matrix $A=I-A_{0}$ designate the contemporaneous causal effects.
-Following physical constraints, we assume that temperature and CO22 do not affect orbital forcing ($a{31}=a_{32}=0$), and that orbital forcing does not directly affect CO22 concentration (a23=0a{23}=0).
-Under these restrictions, Equation (3) yields the following system of structural innovations:
-$\begin{aligned} u_{1t}&=\epsilon_{1t}-a_{12}\epsilon_{2t}-a_{13}\epsilon_{3t}\\ u_{2t}&=-a_{21}\epsilon_{1t}+\epsilon_{2t}\\ u_{3t}&=\epsilon_{3t}\end{aligned}$ (4)
-where $a_{12}$ denotes the Instantaneous Greenhouse Effect (IGE)—the immediate temperature response to a change in CO22 within the same millennium—and a21a{21} denotes the Instantaneous Carbon Feedback (ICF).
-Importantly, $a_{12}=0$ corresponds to a null-IGE identification, where CO2 increases exert no additional warming effect at the contemporaneous resolution.
-Identification of these parameters requires the orthogonality of structural shocks ($\mathbb{E}[u_{it}u_{jt}]=0$ for $i\neq j$).
-Substituting Equation (4) into the condition $\mathbb{E}[u_{1t}u_{2t}]=0$, we derive the Causality Allocation Curve (CAC):
+To isolate structural shocks ($u_{t}$) from the reduced-form residuals ($\epsilon_{t}$), we define the relationship $(I-A_{0})\epsilon_{t}=u_{t}$, where $u_{t}$ is a vector of orthogonal structural innovations with $u_{t}\sim\text{i.i.d.}\mathcal{N}(0,\Sigma_{u})$.
+In our three-variable system, the contemporaneous interaction matrix $A_{0}$ is defined as:
+$A_{0}=\begin{pmatrix}0&a_{12}&a_{13}\\ a_{21}&0&a_{23}\\ 0&0&0\end{pmatrix}$
+where $a_{12}$ and $a_{21}$ represent the contemporaneous temperature response (CTR) and the contemporaneous carbon response (CCR), respectively.
+The third row is set to zero due to the strict exogeneity of orbital forcing, implying $\epsilon_{3t}=u_{3t}$.
+The structural equations for the climate residuals are then given by:
+$u_{1t}$ $=\epsilon_{1t}-a_{12}\epsilon_{2t}-a_{13}\epsilon_{3t}$
+$u_{2t}$ $=-a_{21}\epsilon_{1t}+\epsilon_{2t}-a_{23}\epsilon_{3t}$
+By imposing the orthogonality condition $\mathbb{E}[u_{1t}u_{2t}]=0$ and noting that $\epsilon_{3t}$ is orthogonal to both $u_{1t}$ and $u_{2t}$ (since $\epsilon_{3t}=u_{3t}$), the terms involving orbital forcing effectively vanish in the expectation.
+This yields the Causality Allocation Curve (CAC), which defines the structural trade-off between the CTR and the CCR:
 $(1+a_{12}a_{21})s_{12}-a_{21}s_{11}-a_{12}s_{22}=0$ (2)
-This hyperbola represents all causal allocations $(a_{12},a_{21})$ consistent with the observed residual covariances $s_{ij}$.
+Equation (2) demonstrates that the allocation of the total contemporaneous correlation between temperature and CO2 into directional causal links is uniquely determined by the sample covariance structure ($s_{ij}$) of the residuals.
 Impulse Response Analysis
-To evaluate the dynamic stability and long-term consequences of structural shocks, we employ impulse response functions (IRFs).
-By utilizing the moving average representation of the VECM, the level of the variables at horizon $h$ is expressed as:
-$y_{t+h}=\Psi_{0}u_{t+h}+\Psi_{1}u_{t+h-1}+\dots+\Psi_{h}u_{t}$ (6)
-The matrix $\Psi_{j}$ represents how a structural innovation at time $t$ propagates through the simultaneous interactions and the long-run adjustment mechanism.
-We simulate two hallmark scenarios to quantify climate sensitivity:
+To visualize the propagation of these structural innovations, we utilize Impulse Response Functions (IRFs).
+The IRFs trace the time path of the system in response to a one-time shock in $u_{t}$.
+We simulate two primary scenarios:
   1. 1.
-    A CO2 doubling shock: Represented as a structural innovation in the carbon cycle ($u_{2t}$).
-In our logarithmic specification, a doubling is equivalent to an instantaneous shock of magnitude $\ln(2)\approx 0.693$.
-This allows us to trace the temperature response via the IGE and subsequent feedback loops.
+    A CO2 doubling shock: Represented as $u_{2t}=\ln(2)\approx 0.693$.
+This allows us to observe the transition from the transient response to the long-term Earth System Sensitivity (ESS) determined by $\beta$.
   2. 2.
-    A 2K temperature shock: Represented as an abrupt warming ($u_{1t}=2.0$).
-This scenario quantifies the magnitude of the resulting CO2 release, identifying the strength of the ICF and the efficiency of the error correction process.
-Our VECM is estimated on a 1,000-year grid to ensure a high signal-to-noise ratio in the ice-core proxies.
-While the structural dynamics are defined on this millennial scale, we utilize spline interpolation when presenting the IRFs to facilitate comparison with higher-frequency records without altering the fundamental parameters.
-As derived in Appendix A, the presence of cointegration ensures that temporary structural shocks lead to a permanent shift in the climatic baseline.
-The long-run impact matrix, defined as $\Xi=\lim_{h\to\infty}\Psi_{h}$, represents the total accumulated effect of a shock.
-This property provides a bridge between millennial-scale dynamics and long-term climate transitions, ensuring that the projected trajectories remain robust across diverse time scales.
+    A 2 K temperature shock: Represented as $u_{1t}=2.0$. This isolates the magnitude and speed of the carbon cycle’s response via the CCR and the subsequent feedback loops.
+The presence of cointegration ensures that temporary structural shocks lead to a permanent shift in the climatic baseline.
+The long-run impact matrix, defined as $\Xi=\lim_{h\to\infty}\Psi_{h}$ (see Appendix A, for details), provides a formal bridge between millennial-scale innovations and the long-term equilibrium states.
 Data
-Sources of Paleoclimate and Orbital Data
+Sources
+The empirical foundation of this study rests on three primary data series spanning the late Quaternary: atmospheric CO2 concentrations, Antarctic surface temperature anomalies, and solar insolation.
+Figure 1 illustrates the evolution of these variables over the past 800,000 years, highlighting the strong phase-locking of the climate–carbon system to orbital cycles.
+Atmospheric CO2 data are obtained from the composite ice-core record (Lüthi et al., 2008; Bereiter et al., 2015), which integrates measurements from multiple Antarctic cores to capture natural greenhouse gas variability across eight complete glacial-interglacial cycles.
+For temperature estimates, we utilize the Antarctic temperature anomaly record ($\Delta T$) derived from deuterium ratios in the EPICA Dome C ice core (Jouzel et al., 2007), serving as a reliable indicator for global climatic shifts.
+External orbital forcing is represented by the solar insolation at 60N in June, computed based on the solution by Laskar et al.
+(2004).
+Gridding and Sample Selection
+To conduct a rigorous time-series analysis, the unevenly spaced raw data were transformed into a regular 1,000-year resolution grid.
+We distinguish between two primary sample periods: the comprehensive 800 ka sample and the focused 130 ka sample (covering the period from the Last Interglacial to the present).
+Critically, as visualized in Figure 2, our grid analysis identifies that the 130 ka sample represents a continuous interval entirely free of observational gaps at the millennial resolution.
+Because this interval requires zero interpolation, the resulting series remains a direct reflection of the physical ice-core record, immune to statistical artifacts.
+By focusing on this gap-free 130 ka record, we ensure that the estimated structural parameters—particularly the CCR ($a_{21}$) and structural lag effects—reflect pure physical signals.
 Figure 1: Evolution of climatic and orbital variables over the past 800,000 years.
-The panels display (from top to bottom) solar insolation at $60^{\circ}$N in June and in December, Antarctic temperature anomalies ($\Delta T$), and the logarithm of atmospheric CO2 concentration ($\ln C$), highlighting the strong phase-locking of climate variables to orbital cycles.
-The empirical foundation of this study rests on three primary data series spanning the late Quaternary: atmospheric CO2 concentrations, Antarctic surface temperature anomalies, and solar insolation.
-These datasets provide the necessary temporal coverage to examine the long-term cointegrating relationships and the millennial-scale structural dynamics of the Earth system.
-Atmospheric CO2 data are obtained from the composite ice-core record provided by the World Data Service for Paleoclimatology (Lüthi et al., 2008; Bereiter et al., 2015).
-This high-resolution dataset integrates measurements from the EPICA Dome C, Vostok, and other Antarctic ice cores, extending back to 800,000 years before present (yr BP).
-This record captures the natural variability of greenhouse gases across eight complete glacial-interglacial cycles, providing a robust baseline for evaluating climate sensitivity.
-For temperature estimates, we utilize the Antarctic temperature anomaly record derived from stable isotope (deuterium) ratios in the EPICA Dome C ice core (Jouzel et al., 2007).
-The deuterium record provides a consistent proxy for regional Antarctic temperature, which serves as a reliable indicator for global climatic shifts due to the strong coupling between polar and global temperature variations.
-Solar insolation data, representing the external orbital forcing, are computed using the IMCCE online numerical integration service based on the solution by Laskar et al.
-(2004).
-Specifically, we utilize the mean daily insolation at $60^{\circ}$N during the summer solstice in June.
-This metric is a standard choice (adopted by Lemoine, 2010) for capturing the orbital drivers of glacial-interglacial transitions, commonly referred to as Milankovitch forcing.
-The raw series of these variables, which exhibit clear cyclicality and phase-locking, are plotted in Figure 1.
-Data Gridding and the Selection of Termination II
-The raw ice-core samples are characterized by uneven temporal spacing, which typically becomes sparser with increasing age.
-To conduct a consistent time-series analysis using the Vector Error Correction Model (VECM), the data must be transformed into a regular temporal grid.
-We adopt a two-stage gridding process designed to preserve the integrity of the original physical signal while ensuring statistical reliability.
-In the first stage, the raw data are allocated into 100-year bins.
-When a bin contains multiple observations, the values are averaged to reduce sampling noise and provide a representative value for that centennial interval.
-In the second stage, these 100-year records are aggregated into a 1,000-year grid, which serves as our primary unit of observation ($t$).
-For each 1,000-year grid point, the representative value is calculated by averaging the available 100-year bin values within that specific millennium.
-Crucially, for our primary structural analysis, we do not employ interpolation during this aggregation process.
-If a 1,000-year window contains no underlying 100-year data points, it is intentionally left as a gap.
-This approach ensures that the resulting 1,000-year series remains a direct reflection of the observed ice-core record, free from any statistical artifacts that could be introduced by infilling missing values.
-Linear interpolation at the 100-year level is utilized exclusively for the Whole Period analysis (spanning 800,000 years) to facilitate a continuous estimation over the entire record where data sparsity is more pronounced.
-As illustrated in Figure 2, our grid analysis identifies the presence of observational gaps across the late Quaternary.
-The longest continuous span that is entirely free of such gaps at the 1,000-year resolution—and therefore requires no interpolation whatsoever—perfectly coincides with the period known as Termination II (approximately 141,000 to 121,000 yr BP).
-This period captures the rapid transition from the penultimate glacial maximum to the Eemian interglacial.
-By focusing our primary structural analysis on this gap-free Termination II sample, we ensure that our estimates of the Instantaneous Carbon Feedback (ICF) and the structural lag effects are derived from pure physical signals.
-This rigorous selection process maximizes the internal validity of our causal identification strategy, as it eliminates the possibility that the identified null Instantaneous Greenhouse Effect (IGE) is a product of data smoothing or interpolation bias.
-Figure 2: Visualization of the data gridding structure and observational gaps.
-The upper and middle panels show the density of 100-year and 1,000-year grid points for Temperature and CO2, respectively.
-The bottom panel identifies data gaps (missing observations).
-The period corresponding to Termination II is highlighted as the longest continuous interval free of observational gaps for both variables.
+The panels display (from top to bottom) solar insolation at $60^{\circ}$N in June and December, Antarctic temperature anomalies ($\Delta T$), and the logarithm of atmospheric CO2 concentration ($\ln C$).
+The shaded area indicates the 130 ka sample period used for primary structural analysis.
+Figure 2: Visualization of data gridding and observational gaps.
+The upper panels show the 100-year raw fragments, while the lower panels display the 1,000-year grid used as the VECM basis.
+The dashed box highlights the 130 ka sample as the longest continuous interval free of gaps for both temperature and CO2, justifying its use for structural identification.
 Analysis
@@
 Table 1: Unit Root and Cointegration Test Results
-Termination II (130-130k) Whole Period (800-800k)
+130 ka sample 800 ka sample
 ADF Unit Root Test $Z(t)$ $p$-value $Z(t)$ $p$-value
@@
 Engle-Granger Test $Z(t)$ 10% Crit. $Z(t)$ 1% Crit.
-Cointegration ($T$ on $\ln C$) 3.176-3.176* 3.077-3.077 7.871-7.871*** 3.910-3.910
+Cointegration ($T$ on $\ln C$) 3.176-3.176 3.077-3.077 7.871-7.871∗∗∗ 3.910-3.910
 Notes: ADF tests are performed with a constant.
@@
 Table 2: Lag Order Selection Criteria
-Termination II Whole Period
+130 ka sample 800 ka sample
 Lag FPE AIC HQIC SBIC FPE AIC HQIC SBIC
@@
 2 $0.0002$ 3.047-3.047 2.957-2.957 2.824-2.824 $0.0001$ 3.301-3.301 3.278-3.278 3.242-3.242
-3 0.00010.0001* 3.157-3.157* 3.030-3.030* 2.845-2.845* $0.0001$ 3.375-3.375 3.343-3.343* 3.292-3.292*
-4 $0.0002$ 3.108-3.108 2.945-2.945 2.707-2.707 0.00010.0001* 3.375-3.375* 3.334-3.334 3.269-3.269
+3 $0.0001$ 3.157-3.157 3.030-3.030 2.845-2.845 $0.0001$ 3.375-3.375 3.343-3.343 3.292-3.292
+4 $0.0002$ 3.108-3.108 2.945-2.945 2.707-2.707 $0.0001$ 3.375-3.375 3.334-3.334 3.269-3.269
 Notes: * indicates lag order selected by the criterion. AIC: Akaike info criterion (FPE: final prediction error); SBIC: Schwarz Bayesian info criterion; HQIC: Hannan-Quinn info criterion.
-Table 3: Comparison of VECM Results: Termination II vs. Whole Period
-Termination II Whole Period
+Table 3: Comparison of VECM Results: 130 ka vs. 800 ka samples
+130 ka sample 800 ka sample
 Panel A: Cointegrating Equation ($\beta$)
@@
 $T_{t-1}$ $1.000$ $1.000$
-$\ln C_{t-1}$ 18.748-18.748*** $(2.312)$ 20.251-20.251*** $(1.081)$
-$M_{t-1}$ $0.009$ $(0.038)$ 0.105-0.105*** $(0.017)$
+$\ln C_{t-1}$ 18.748-18.748∗∗∗ $(2.312)$ 20.251-20.251∗∗∗ $(1.081)$
+$M_{t-1}$ $0.009$ $(0.038)$ 0.105-0.105∗∗∗ $(0.017)$
 Constant $103.300$ $164.974$
@@
 Independent Variables $\Delta T_{t}$ $\Delta\ln C_{t}$ $\Delta T_{t}$ $\Delta\ln C_{t}$
-Error Correction ($\alpha$) 0.103-0.103** $0.002$ 0.164-0.164*** 0.0002-0.0002
+Error Correction ($\alpha$) 0.103-0.103∗∗ $0.002$ 0.164-0.164∗∗∗ 0.0002-0.0002
 $(0.042)$ $(0.001)$ $(0.017)$ $(0.001)$
-$\Delta T_{t-1}$ $0.030$ $0.003$ 0.3460.346*** 0.0030.003***
+$\Delta T_{t-1}$ $0.030$ $0.003$ 0.3460.346∗∗∗ 0.0030.003∗∗∗
 $(0.096)$ $(0.003)$ $(0.034)$ $(0.001)$
-$\Delta T_{t-2}$ 0.281-0.281*** $0.0004$ 0.123-0.123*** 0.0030.003**
+$\Delta T_{t-2}$ 0.281-0.281∗∗∗ $0.0004$ 0.123-0.123∗∗∗ 0.0030.003∗∗
 $(0.093)$ $(0.003)$ $(0.033)$ $(0.001)$
-$\Delta\ln C_{t-1}$ 15.20115.201*** 0.4470.447*** 6.0256.025*** 0.6000.600***
+$\Delta\ln C_{t-1}$ 15.20115.201∗∗∗ 0.4470.447∗∗∗ 6.0256.025∗∗∗ 0.6000.600∗∗∗
 $(2.845)$ $(0.102)$ $(1.033)$ $(0.035)$
-$\Delta\ln C_{t-2}$ 0.514-0.514 0.307-0.307*** $0.723$ 0.312-0.312***
+$\Delta\ln C_{t-2}$ 0.514-0.514 0.307-0.307∗∗∗ $0.723$ 0.312-0.312∗∗∗
 $(3.008)$ $(0.108)$ $(1.080)$ $(0.037)$
-$\Delta M_{t-1}$ 0.093-0.093*** 0.0005-0.0005 0.1460.146*** $0.0000$
+$\Delta M_{t-1}$ 0.093-0.093∗∗∗ 0.0005-0.0005 0.1460.146∗∗∗ $0.0000$
 $(0.033)$ $(0.001)$ $(0.020)$ $(0.001)$
@@
 Table 1 reports the results of the Augmented Dickey-Fuller (ADF) unit root tests and the Engle-Granger cointegration tests.
-The ADF statistics indicate that both temperature and the logarithm of CO2 are non-stationary $I(1)$ processes across both the Whole Period and the Termination II samples.
+The ADF statistics indicate that both temperature and the logarithm of CO2 are non-stationary $I(1)$ processes across both the 800 ka and the 130 ka samples.
 The Engle-Granger tests confirm the existence of a stable cointegrating relationship, providing the empirical basis for the VECM specification defined in Equation (1).
 Furthermore, the lag order selection criteria presented in Table 2 suggest that a second-order lag structure (where the level-VAR lag $k=3$ corresponds to VECM lag $p=2$) is optimal for capturing the millennial-scale dynamics while maintaining statistical parsimony.
-The primary estimation results for the VECM are summarized in Table 3.
-Panel A presents the identified cointegrating vector ($\beta$), which defines the long-run physical bond between the variables.
-In both samples, we find a remarkably consistent and statistically significant equilibrium relationship.
-Specifically, for the Termination II sample—which provides the most reliable signal due to the absence of interpolation gaps—the relationship is estimated as $T-18.748\ln C=\textit{const}$.
-From this result, we derive the Earth System Sensitivity (ESS), defined as the total temperature response to a doubling of atmospheric CO2 concentration after all long-term feedbacks have reached equilibrium:
-$\text{ESS}=18.748\times\ln(2)\approx 13.0\text{ K}$ (7)
-Our estimate of 13.0 K is situated at the higher end of the range reported in the literature, such as the $9\pm 4$ K (95% CI: 7–13 K) estimate by Snyder (2016).
-This high sensitivity can be understood through the conceptual framework of the PALAEOSENS project (Rohling et al., 2012).
-While the conventional Equilibrium Climate Sensitivity (ECS) typically accounts for “fast” feedbacks such as water vapor and clouds (often referred to as Charney sensitivity), our VECM, calibrated on the millennial-scale transitions of the Pleistocene, inherently captures the “slow” feedbacks of the Earth system.
-These include ice-sheet albedo changes, vegetation shifts, and long-term carbon-cycle responses that operate over multi-millennial time scales.
-According to Rohling et al.
-(2012), the inclusion of these slow feedbacks—which characterizes the transition from $S_{act}$ to $S_{ESS}$—can significantly amplify the total temperature response, potentially doubling the sensitivity compared to the Charney estimate.
-Our result of 13.0 K represents this comprehensive Earth System Sensitivity ($S_{ESS}$), reflecting the state-dependent amplification characteristic of major deglaciation events like Termination II.
-Panel B of Table 3 delineates the short-run dynamics and the adjustment mechanisms of the system.
-The reduced-form lag coefficient for CO2 in the temperature equation, $(\Gamma_{1})_{12}=15.201$ ($p<0.01$), identifies the impact of CO2 changes from the previous millennium.
-Crucially, under the assumption of a null Instantaneous Greenhouse Effect ($a_{12}=0$)—a condition to be empirically validated in the subsequent section via the ICF Window for null IGE—this reduced-form parameter directly identifies the structural lag coefficient $(\Gamma_{1}^{*})_{12}$, representing the momentum of the greenhouse effect.
-By recalling the structural identity $(I-A_{0})\Delta y_{t}=(I-A_{0})\alpha\beta^{\prime}y_{t-1}+\Gamma^{*}_{1}\Delta y_{t-1}+\dots+u_{t}$, the relationship between the structural and reduced-form coefficients is defined by the identity $\Gamma^{*}_{1}=(I-A_{0})\Gamma_{1}$.
-The specific element for the CO2-to-temperature lag can be expanded as follows:
-$(\Gamma^{*}_{1})_{12}=(\Gamma_{1})_{12}-a_{12}(\Gamma_{1})_{22}-a_{13}(\Gamma_{1})_{32}$ (8)
-Owing to the strict exogeneity of orbital parameters, past climate states cannot influence current orbital forcing, which dictates that the reduced-form coefficient $(\Gamma_{1})_{32}$ is effectively zero.
-Under this physical constraint, and imposing the null-IGE identification ($a_{12}=0$), the second and third terms in the expansion vanish.
-Consequently, the structural lag effect is directly identified by the estimated reduced-form coefficient: $(\Gamma^{*}_{1})_{12}=(\Gamma_{1})_{12}=15.201$.
-This structural identification allows for a precise decomposition of the total temperature response observed in the 1,000-year window following a CO2 doubling event ($\Delta\ln C=\ln 2\approx 0.693$).
-The resulting warming is the sum of two distinct physical forces:
-  1.
-Structural Momentum: The structural lag effect $(\Gamma^{*}_{1})_{12}$ provides an immediate "kick" of $15.201\times\ln(2)\approx 10.5$ K.
-This represents the powerful, delayed radiative feedback that manifests after the initial 1,000-year period of relative stability defined by the null-IGE identification.
-  2.
-Equilibrium Adjustment: The error correction coefficient $\alpha=-0.103$ ($p=0.013$) quantifies the system’s restoring force toward the long-run ESS.
-Starting from the disequilibrium created by the CO2 innovation, this mechanism contributes an additional adjustment of approximately $13.0\text{ K}\times 0.103\approx 1.3$ K.
-Combined, these forces generate a total warming of approximately 11.8 K within the first millennium after the lag period.
-The fact that this combination of delayed structural momentum and equilibrium pull so effectively accounts for the millennial-scale temperature variance explains why the model identifies $a_{12}\approx 0$ at the contemporaneous scale.
-In our framework, the greenhouse effect is not diminished; rather, it is structurally localized as a delayed feedback, allowing the contemporaneous driver of climatic shifts to be identified as orbital-induced temperature changes.
-4.1.1 Consistency with Modern Anthropogenic Warming
-The robustness of our SVECM estimates is further validated by comparing the model’s projected warming rate with observed modern temperature trends.
-Our results for Termination II indicate that a doubling of CO2 generates a total millennial response of approximately 11.8 K, comprising the 10.5 K structural momentum and the 1.3 K equilibrium adjustment.
-This translates to a centennial warming rate of 1.18 K per 100 years for a CO2 doubling.
-To contextualize this within the modern era, we consider the increase in atmospheric CO2 since the pre-industrial period, from a baseline of approximately 280 ppm to the current level of roughly 428 ppm.
-In logarithmic terms, this corresponds to an accumulation of approximately 61% of a doubling:
-$\text{Modern Accumulation Factor}=\frac{\ln(428/280)}{\ln(2)}\approx 0.61$ (9)
-Applying our derived centennial rate to this factor, the model predicts a CO2-induced warming of:
-$1.18\text{ K/century}\times 0.61\approx 0.72\text{ K/century}$ (10)
-Remarkably, this paleo-derived estimate is in near-perfect alignment with the observed instrumental warming rate of approximately 0.73–0.80 K per century (equivalent to the observed 1.1–1.2 K rise over the last 150 years).
-This high degree of correspondence suggests that the null-IGE identification ($a_{12}=0$) at the millennial scale does not diminish the long-term radiative importance of CO2.
-Instead, it reveals that the intrinsic sensitivity of the Earth system—as captured during the rapid transitions of the Pleistocene—is sufficient to account for the bulk of modern warming through delayed structural feedbacks and equilibrium adjustments.
-The fact that a model calibrated on 130,000-year-old ice-core data can so accurately recover the scale of 21st-century climate change provides powerful empirical support for our causal identification strategy.
+The primary estimation results for the VECM, comparing the gap-free 130 ka sample with the interpolated 800 ka whole period, are summarized in Table 3.
+Panel A presents the estimated cointegrating vector ($\beta$), which defines the long-term equilibrium relationship between temperature and CO2.
+In both samples, the coefficient for $\ln C_{t-1}$ is negative and highly significant ($p<0.01$), indicating a robust positive long-term correlation between climatic warming and atmospheric carbon concentration.
+Panel B describes the short-run dynamics and the error correction mechanism.
+The error correction term ($\alpha$) for the temperature equation is negative and statistically significant in both samples (-0.103 for the 130 ka sample), confirming that temperature deviations from the long-term carbon-linked equilibrium are corrected over time.
+In contrast, the error correction term for the $\ln C_{t}$ equation is statistically indistinguishable from zero, suggesting that atmospheric CO2 acts as the primary stochastic driver of the long-term trend in this system.
+The short-run lagged effects ($\Delta T_{t-i},\Delta\ln C_{t-i}$) also reveal significant millennial-scale persistence and cross-variable interactions.
+These estimated parameters provide the necessary foundation for the structural causality analysis and sensitivity evaluations performed in the subsequent sections.
 Causality Allocation Analysis
-4.2.1 Instantaneous Carbon Feedback (ICF)
 The structural parameter $a_{21}$ is defined as a key element of the contemporaneous interaction matrix $A_{0}$ in our SVAR system, representing the simultaneous feedback within a single 1,000-year time step.
-Specifically, $a_{21}$ quantifies the Instantaneous Carbon Feedback (ICF), defined as the response of the logarithm of CO2 concentration to a change in temperature:
-$a_{21}=\frac{\partial\ln C_{t}}{\partial T_{t}}$ (11)
+Specifically, $a_{21}$ quantifies the contemporaneous carbon response (CCR), defined as the response of the logarithm of CO2 concentration to a change in temperature:
+$a_{21}=\frac{\partial\ln C_{t}}{\partial T_{t}}$
 To estimate this parameter from the paleoclimate record, we utilize a regression model formulated in first differences:
@@
 Given the simultaneous nature of the coupling, a reverse causal link also exists:
-$\Delta T_{t}=a_{12}\Delta\ln C_{t}+w_{t}$ (13)
-where $a_{12}$ represents the Instantaneous Greenhouse Effect (IGE).
-If a significant IGE exists at the millennial resolution, the OLS estimate for $a_{21}$ would suffer from simultaneity bias, as $\Delta T_{t}$ would be correlated with the error term $v_{t}$.
+$\Delta T_{t}=a_{12}\Delta\ln C_{t}+w_{t}$
+where $a_{12}$ represents the contemporaneous temperature response (CTR).
+If a significant CTR exists at the millennial resolution, the OLS estimate for $a_{21}$ would suffer from simultaneity bias, as $\Delta T_{t}$ would be correlated with the error term $v_{t}$.
 To address this concern and ensure the consistency of our estimates, we first employ Instrumental Variables (IV) estimation.
@@
 The estimation results are summarized in Table 4.
-Crucially, the endogeneity test (a robust Durbin-Wu-Hausman test) for the Termination II sample yields a $p$-value of 0.864.
+Crucially, the endogeneity test (a robust Durbin-Wu-Hausman test) for the 130 ka sample yields a $p$-value of 0.864.
 This result indicates that we cannot reject the null hypothesis that the temperature change is exogenous with respect to CO2 innovations.
-In the context of our causality allocation, this provides strong empirical evidence that the contemporaneous reverse causality from CO2 to temperature (the IGE) is statistically undetectable at the 1,000-year resolution.
+In the context of our causality allocation, this provides strong empirical evidence that the contemporaneous reverse causality from CO2 to temperature (the CTR) is statistically undetectable at the 1,000-year resolution.
 Consequently, OLS estimation is not only consistent but also more efficient than IV estimation.
-The resulting OLS estimate for Termination II is $\hat{a}_{21}=0.02082$ ($z=8.04,p<0.01$), which we adopt as the representative structural parameter for the ICF.
-Table 4: IV Estimation Results for the Instantaneous Carbon Feedback ($a_{21}$)
+The resulting OLS estimate is $\hat{a}_{21}=0.02082$ ($z=8.04,p<0.01$), which we adopt as the representative structural parameter for the CCR.
+Table 4: IV Estimation Results for CCR ($a_{21}$)
 Sample Period Estimate $z$ 95% Conf. Interval Sensitivity
-$($$a_{21}$) Lower Upper (ppm/K)
-Whole Period $0.0222$ $5.44$ $0.0142$ $0.0302$ $6.22$
-Termination II $0.0219$ $3.81$ $0.0106$ $0.0332$ $6.13$
-Termination II (OLS) $0.02082$ $8.04$ $0.0157$ $0.0259$ $5.82$
+$a_{21}$ Lower Upper (ppm/K)
+800 ka $0.0222$ $5.44$ $0.0142$ $0.0302$ $6.22$
+130 ka $0.0219$ $3.81$ $0.0106$ $0.0332$ $6.13$
+130 ka (OLS) $0.0208$ $8.04$ $0.0157$ $0.0259$ $5.82$
 Diagnostics for IV Estimation 1st $F$ Overid. Endog. Obs.
-Whole Period $28.48$ $0.963$ $0.005$ 801
-Termination II $11.91$ $0.054$ $0.864$ 130
+800 ka $28.48$ $0.963$ $0.005$ 801
+130 ka $11.91$ $0.054$ $0.864$ 130
 Notes: The dependent variable is $\Delta\ln C_{t}$ and the endogenous regressor is $\Delta T_{t}$ as defined in Equation (3).
-Instruments for the Whole Period are $\{M^{\text{N60J}}_{t},M^{\text{N60D}}_{t}\}$ and for Termination II are $\{M^{\text{N60D}}_{t},T_{t-2}\}$.
+Instruments for 800 ka are $\{M^{\text{N60J}}_{t},M^{\text{N60D}}_{t}\}$ and for 130 ka are $\{M^{\text{N60D}}_{t},T_{t-2}\}$.
 Overid.
@@
 Endog.
-reports the $p$-value for the Durbin-Wu-Hausman endogeneity test; the high $p$-value for Termination II (0.864) indicates that the null hypothesis of exogeneity for $\Delta T_{t}$ cannot be rejected, justifying the use of the OLS estimate ($\hat{a}_{21}=0.02082$).
-The convergence between OLS and IV estimates supports the identification of a null instantaneous greenhouse effect ($a_{12}=0$).
+reports the $p$-value for the Durbin-Wu-Hausman endogeneity test; the high $p$-value 130 ka (0.864) indicates that the null hypothesis of exogeneity for $\Delta T_{t}$ cannot be rejected, justifying the use of the OLS estimate ($\hat{a}_{21}=0.0208$).
 Sensitivity is converted as $a_{21}\times 280$ ppm/K for comparison.
-4.2.2 Instantaneous Greenhouse Effect (IGE)
-The absence of a significant contemporaneous IGE ($a_{12}=0$), as suggested by the endogeneity diagnostics, can be further explored and validated through the Causality Allocation Curve (CAC).
-This curve represents the set of ($a_{12},a_{21}$) pairs that are consistent with the estimated reduced-form covariance matrix of the VECM residuals, assuming the orthogonality of structural innovations.
-As illustrated in Figure 3, the CAC defines the structural trade-off between the allocated strengths of IGE and ICF.
-A remarkable consistency is observed between two independent analytical paths.
-The ICF value identified by our first-difference OLS ($\hat{a}_{21}=0.02082$) almost perfectly coincides with the structural $a_{21}$ value where the CAC intersects the $a_{12}=0$ axis ($\hat{a}_{21}=0.02084$), which was estimated independently via the Method of Moments based on the system-wide VECM residuals (as reported in Table 5).
-This convergence between instrumental variable diagnostics and system-wide residual analysis is a significant finding.
-It suggests that the model, including the VECM lag structure and the null-IGE identification ($a_{12}=0$), captures the underlying physical dynamics of the Pleistocene with high fidelity.
-If the model had been mis-specified, or if a significant instantaneous greenhouse effect were truly present, the OLS estimate would be expected to diverge significantly from the point where $a_{12}=0$.
-To quantify the statistical stability of this identification, we define the ICF Window for null IGE as the range of ICF values for which the null hypothesis of $a_{12}=0$ cannot be rejected.
-As shown in Panel B of Table 5, a sensitivity analysis conducted at the boundaries of the 95% confidence interval for our ICF estimate confirms that the IGE remains statistically indistinguishable from zero across the entire plausible range of carbon cycle feedback.
-Furthermore, we consider the median sensitivity estimate from Frank et al.
-(2010), which corresponds to $a_{21}\approx 0.0275$ when converted to our logarithmic scale (7.7 ppm/K divided by a 280 ppm baseline).
-While this value sits near the boundary of the ICF Window for null IGE (being rejected at the 5% level but not at the 1% level), our data-driven estimate of 0.0208 lies precisely at its center.
-This indicates that while the consensus in the carbon-cycle literature is broadly compatible with a temperature-led causal structure, our empirical identification achieves a state of maximal statistical consistency.
-These results collectively reinforce the conclusion that the greenhouse effect in the paleoclimate is not an instantaneous phenomenon but is instead manifested through the structural lagged effects and long-run equilibrium adjustments described in the previous section.
-Table 5: $M$ Estimation Results for ICF Window
-Restriction / Parameter Estimate $z$ $P>|z|$ 95% Conf. Interval
-Lower Upper
-Panel A: Estimating $a_{21}$ assuming null marginal greenhouse forcing ($a_{12}=0$)
-$a_{21}$ (Carbon Feedback) $0.02084$ $7.81$ $0.000$ $0.0156$ $0.0261$
-Panel B: Sensitivity analysis of $a_{12}$ at the boundaries of the $a_{21}$ 95% CI
-$a_{12}$ (at $a_{21}=0.0156$) $5.3805$ $2.00$ $0.046$ $0.1052$ $10.6557$
-$a_{12}$ (at $a_{21}=0.0261$) 6.9366-6.9366 1.71-1.71 $0.087$ 14.8738-14.8738 $1.0006$
-Notes: Estimates are obtained using the Method of Moments (MM) based on VECM residuals.
-Panel A identifies the ICF value required to satisfy the condition of a null instantaneous greenhouse effect ($a_{12}=0$).
-Panel B demonstrates the statistical stability of the IGE estimate at the boundaries of the ICF’s 95% confidence interval.
-The resulting ICF Window for null IGE encompasses our empirical OLS estimate, confirming the structural consistency of the model.
-Figure 3: The Causality Allocation Curve and ICF Window for null IGE.
-The curve illustrates the structural trade-off between the Instantaneous Greenhouse Effect (IGE, $a_{12}$) and the Instantaneous Carbon Feedback (ICF, $a_{21}$).
-The white region identifies the ICF Window for null IGE, where the hypothesis $a_{12}=0$ is not rejected at the 1% level.
-Our OLS estimate for Termination II ($\hat{a}_{21}=0.02082$) lies centrally within this window, highlighting the robustness of the null-IGE identification.
-Impulse Response Simulation of Climate-Carbon Dynamics
-In this final part of the analysis, we evaluate the dynamic stability and the long-term sensitivity of the identified Earth system via impulse response functions (IRFs).
-By utilizing the moving average representation of the estimated SVECM, we simulate the propagation of structural innovations across millennial time scales.
-These simulations provide a visual and quantitative synthesis of the Instantaneous Carbon Feedback (ICF), the Instantaneous Greenhouse Effect (IGE), and the structural lagged effects identified in the preceding sections.
-Crucially, the IRF trajectories allow us to distinguish between the immediate structural impacts and the long-term adjustment toward the ESS discussed in Section 4.1.
-As we move from the instantaneous response (governed by $a_{ij}$) toward the multi-millennial equilibrium, the transition captures the progression from the initial radiative forcing to the fully realized Earth System Sensitivity.
-This dynamic visualization aligns with the perspective of Rohling et al.
-(2012), who emphasize that the measured sensitivity is a function of the time scale considered, with the full magnitude of slow feedbacks only becoming apparent as the system approaches its new steady state.
-4.3.1 System Response to a CO2 Doubling Innovation
-We first simulate a scenario representing an abrupt doubling of atmospheric CO2 concentration.
-This is implemented as a structural innovation in the carbon cycle, $u_{2t}=\ln(2)\approx 0.693$, applied to a system initially at a stable equilibrium (baseline CO2 = 280 ppm).
-The resulting trajectories for temperature and CO2 are presented in Figure 4.
-As shown in the left panels of Figure 4, the model captures a distinct two-stage temperature response.
-During the first 1,000-year interval immediately following the shock, the temperature anomaly remains effectively zero.
-This period of stasis is a direct consequence of our null-IGE identification ($a_{12}=0$), which implies that CO2 concentrations do not exert an additional thermal impact within the same millennial time step.
-However, as the system enters the second millennium, we observe a sharp and powerful rise in temperature.
-This "thermal breakout" is driven by the structural lag coefficient $(\Gamma^{*}_{1})_{12}=15.201$, which delivers a structural momentum of approximately 10.5 K.
-Simultaneously, the error correction mechanism ($\alpha=-0.103$) begins to pull the system toward its long-run Earth System Sensitivity (ESS).
-Over the subsequent 10,000 years, the temperature continues to rise as the slow-acting feedbacks of the Pleistocene climate reach maturity.
-As indicated by the long-run impact matrix $\Xi$, the temperature eventually stabilizes at a permanent increase of approximately 13.0 K, consistent with the ESS derived from the cointegrating vector.
-This simulation illustrates that the greenhouse effect in the paleoclimate is not an instantaneous force but a structurally delayed phenomenon.
-The ability of the model to account for the observed variance in temperature through this delayed momentum, while maintaining $a_{12}=0$, provides a compelling physical explanation for the phase-locking observed in the ice-core records.
+To further rigorously test the statistical significance of the CTR ($a_{12}$) while accounting for the uncertainty in the CCR ($a_{21}$), we perform a joint estimation using the Generalized Method of Moments (GMM).
+This approach incorporates both the regression model in Equation (3) and the orthogonality condition derived from the structural shock properties, known as the Causality Allocation Curve (CAC) or Equation (2).
+The moment conditions for this joint estimation are defined by the structural identification framework established in Section 2.2.
+The results of the GMM joint estimation are presented in Table 5.
+The estimate for $a_{21}$ is 0.0206 ($z=7.24,p<0.01$), which is highly consistent with the earlier OLS result.
+In contrast, the estimate for $a_{12}$ is 0.2209 with an exceptionally large standard error of 2.5253.
+A Wald test for the null hypothesis $H_{0}:a_{12}=0$ yields a $\chi^{2}(1)$ statistic of 0.01 with a $p$-value of 0.9303, meaning we cannot reject the hypothesis that the CTR is zero.
+This robustly confirms that, at the millennial resolution, the climate-carbon feedback is characterized by a unidirectional response from temperature to carbon, without any detectable instantaneous reverse causality.
+Table 5: GMM Joint Estimation Results and Wald Test for Structural Parameters
+Coef. Std. Err. 95% Conf. Interval Sensitivity
+Parameter Lower Upper (ppm/K)
+$a_{21}$ 0.0206∗∗∗ $0.0029$ $0.0151$ $0.0262$ $5.77$
+$a_{12}$ $0.2209$ $2.5253$ 4.7286-4.7286 $5.1703$
+Constant $0.0011$ $0.0018$ 0.0024-0.0024 $0.0046$
+Wald Test of $H_{0}:a_{12}=0$
+$\chi^{2}(1)=0.01\qquad\text{Prob}>\chi^{2}=0.9303$
+Notes: Sensitivity is converted as $a_{21}\times 280$ ppm/K for comparison. ∗∗∗ denotes statistical significance at the 1% level.
+Figure 3 visualizes this identification.
+The plot displays the theoretical Causality Allocation Curve (CAC), representing the locus of $(a_{21},a_{12})$ pairs satisfying the shock orthogonality condition.
+The GMM estimate and its 95% confidence ellipse are centered near the $a_{12}=0$ axis.
+The horizontal elongation of the ellipse reflects the lack of empirical evidence for a non-zero CTR, while the tight interval for the CCR ($a_{21}$) emphasizes its stability across estimation methods.
+Figure 3: Identification of structural parameters on the Causality Allocation Curve (CAC).
+The blue line represents the CAC derived from the residual covariance.
+The red point and dashed ellipse indicate the GMM joint estimate and the 95% confidence region.
+The intersection near the $a_{12}=0$ line confirms the unidirectional contemporaneous feedback from temperature to carbon.
+Structural Sensitivity and Response
+Based on the identification results in the previous section, we now evaluate the structural sensitivity of the climate–carbon system.
+First, to interpret the estimated reduced-form parameters ($\alpha,\beta,\Gamma$) in a structural context, we must consider the transformation $(I-A_{0})$.
+While the cointegrating vector $\beta$ and the adjustment speeds $\alpha$ remain invariant to this transformation due to the property $\beta^{\prime}\Xi=0$ (as formally derived in Appendix A), the short-run lag coefficients $\Gamma$ must be structurally re-identified using the contemporaneous parameters $a_{12}$ and $a_{21}$.
+By recalling the structural identity $(I-A_{0})\Delta y_{t}=(I-A_{0})\alpha\beta^{\prime}y_{t-1}+\Gamma^{*}_{1}\Delta y_{t-1}+\dots+u_{t}$, the relationship between the structural and reduced-form coefficients is defined by the identity $\Gamma^{*}_{1}=(I-A_{0})\Gamma_{1}$.
+The specific element for the CO22-to-temperature lag can be expanded as follows:
+$(\Gamma^{*}_{1})_{12}=(\Gamma_{1})_{12}-a_{12}(\Gamma_{1})_{22}-a_{13}(\Gamma_{1})_{32}$ (4)
+Owing to the strict exogeneity of orbital parameters, past climate states cannot influence current orbital forcing, which dictates that the reduced-form coefficient $(\Gamma_{1})_{32}$ is effectively zero.
+Under this physical constraint, and imposing the null-CTR identification ($a_{12}=0$) established by our GMM results, the second and third terms in the expansion in Equation (4) vanish.
+Consequently, the structural lag effect is directly identified by the estimated reduced-form coefficient: $(\Gamma^{*}_{1})_{12}=(\Gamma_{1})_{12}=15.201$.
+This structural consistency allows us to derive the Earth System Sensitivity (ESS) directly from the cointegrating relationship.
+The estimated long-term equilibrium for the 130 ka sample is given by $T-18.748\ln C=\text{const}$.
+This relationship implies that a doubling of atmospheric CO2 concentration is associated with a long-term temperature increase of $18.748\times\ln(2)\approx 13.0$ K in Antarctica.
+However, since this estimate is based on Antarctic ice core records, it reflects local temperature variations which are subject to polar amplification.
+Following the methodology of previous studies, we convert this local signal to a Global Average Temperature (GAT) equivalent by applying a scaling factor of approximately 0.5, as Antarctic temperature anomalies are known to be roughly twice as large as the global mean.
+This conversion yields a global ESS of approximately 6.5 K per CO2 doubling.
+This value is remarkably consistent with the reconstructed ESS of $9\pm 4.5$ K (95% CI) reported by Snyder (2016), confirming that our VECM-based structural identification captures the fundamental long-term coupling of the Earth’s climate and carbon cycle since the Last Interglacial.
+Beyond the long-term equilibrium, the structural VECM allows for a granular decomposition of the transient temperature response.
+We examine the impact of a doubling of atmospheric CO22 concentration ($\Delta\ln C=\ln(2)$) within a single millennial time step.
+Under the structural identification a12=0a{12}=0, the immediate temperature response after 1,000 years is driven by two distinct mechanisms: the structural lag effect and the error correction process.
+First, the structural lag effect $(\Gamma^{*}_{1})_{12}=15.201$ accounts for a temperature increase of $15.201\times\ln(2)\approx 10.5$ K.
+This represents the direct millennial-scale sensitivity of the climate system to a carbon shock.
+Second, the error correction term $\alpha=-0.103$ dictates the speed at which the system reverts to its long-term equilibrium.
+A doubling of CO2 creates a disequilibrium in the cointegrating relationship of $18.748\times\ln(2)\approx-13.0$ K.
+The resulting adjustment in the subsequent millennium contributes an additional warming of $(-0.103)\times(-13.0)\approx 1.3$ K.
+Summing these effects, the cumulative temperature response within one millennium of a CO2 doubling reaches $11.8$ K in Antarctica.
+In global average terms, this corresponds to approximately $5.9$ K of warming.
+While this magnitude appears large, it is remarkably consistent with the transient warming observed in the recent industrial era.
+For instance, the observed global temperature increase of approximately $1.2$ K over the last 150 years, associated with a CO2 rise from 280 to 428 ppm (a 53% increase), aligns with the sensitivity scale identified in our millennial-resolution model when accounting for the slower response times of the Earth system.
+Summing these effects, the cumulative temperature response within one millennium of a CO2 doubling reaches 11.8 K in Antarctica, which translates to a global average warming of approximately 5.9 K.
+While this 1,000-year response is substantial, it represents only the transient stage of the total adjustment toward the long-term equilibrium (ESS) of 6.5 K.
+The proximity of our VECM-based estimates to the results reported by Snyder (2016) warrants further methodological discussion.
+Snyder (2016) estimated the ESS using a simple OLS framework and was subsequently criticized for potentially ignoring the endogeneity of CO2 and the influence of orbital forcing.
+However, our identification of a stable cointegrating relationship ($\beta$) provides a strong theoretical justification for the robustness of such OLS-based reconstructions.
+In the presence of cointegration, OLS estimators possess the property of “super-consistency,” converging to the true long-run parameters even in the face of contemporaneous endogeneity.
+By employing a VECM, we have formally verified this long-term stability and separated the structural feedbacks from stochastic innovations.
+The fact that our structurally identified ESS (6.5 K) and millennial response (5.9 K) align closely with Snyder’s findings suggests that the climate–carbon coupling is so fundamentally dominated by the long-term cointegrating trend that the essence of the Earth system’s sensitivity is captured even by simpler models, provided they span sufficient glacial cycles.
+Our structural approach thus reinforces the empirical validity of previous reconstructions while providing a rigorous framework to account for internal dynamics and orbital exogeneity.
+This millennial sensitivity also provides a critical perspective on modern climate change.
+A global warming of 5.9 K per CO2 doubling after 1,000 years implies an average linear adjustment rate of approximately 0.885 K per 150 years.
+While this is numerically lower than the observed 1.2 K increase since the industrial revolution (associated with a 53% rise in CO2), this discrepancy is entirely consistent with the non-linear “front-loading” characteristic of climate responses.
+The immediate radiative forcing following a carbon shock triggers rapid warming in the upper ocean and atmosphere (often termed the Transient Climate Response), whereas our 6.5 K ESS estimate captures the slow, “committed” warming of the deep Earth system that unfolds over subsequent centuries.
+Rather than a contradiction, the gap between the modern 1.2 K response and our 1,000-year 5.9 K projection quantifies the massive, yet-to-be-realized thermal debt inherent in the Earth system’s long-term feedbacks.
+Finally, we evaluate our estimate of the Contemporaneous Carbon Response (CCR).
+Our GMM estimation yields $\hat{a}_{21}=0.0206$, which corresponds to a carbon cycle sensitivity of approximately $5.77$ ppm per 1 K of Antarctic warming.
+This result is highly consistent with the median estimate of $7.7$ ppm/K (likely range: 1.7–21.4 ppm/K) reported by Frank et al.
+(2010) for the last millennium.
+The stability of the CCR across different timescales—from the centennial resolution of the late Holocene to the millennial resolution of our 130 ka sample—underscores the robustness of the climate–carbon feedback architecture identified in this study.
+Impulse Response Simulation of Climate–Carbon Dynamics
+To visually clarify the feedback loops identified in our structural VECM, we perform impulse response function (IRF) simulations for two hallmark scenarios: a CO2 doubling shock ($u_{2t}=\ln 2$) and a sudden 2 K temperature increase ($u_{1t}=2.0$).
+Figure 4 illustrates the response to the CO2 doubling shock.
+Consistent with our null-CTR identification ($a_{12}=0$), temperature shows no immediate response within the same 1,000-year step.
+Instead, the warming is triggered by structural lag effects, gradually converging toward the long-term ESS of 13.0 K in Antarctica.
+This dynamic visualization aligns with Rohling et al.
+(2012), who emphasize that climate sensitivity is a function of the timescale, with the full magnitude of slow feedbacks becoming apparent only as the system approaches a new steady state.
+In contrast, Figure 5 displays the response to a 2 K temperature shock.
+Here, the carbon cycle responds instantaneously, reflecting the statistically significant CCR ($a_{21}>0$).
+This immediate CO2 release further amplifies the temperature anomaly through reciprocal lag effects, characterizing the self-reinforcing nature of the climate–carbon system.
+The distinct contrast between these two simulations provides clear visual evidence of the unidirectional contemporaneous causality at millennial resolutions.
 Figure 4: Impulse response functions for a CO2 doubling shock ($u_{2t}=\ln 2$).
-The plots show the evolution of CO2 concentration (upper panels) and Antarctic temperature anomalies (lower panels) over a 100,000-year horizon (left) and a detailed 10,000-year view (right).
-Note the 1,000-year lag in the temperature response, followed by a transition toward the long-run ESS of 13.0 K.
+The plots show the evolution of CO2 concentration (upper) and Antarctic temperature (lower) over 100,000 years (left) and 10,000 years (right).
+The 1,000-year lag in temperature confirms $a_{12}=0$, while the gradual rise illustrates the time-dependent nature of ESS.
 Figure 5: Impulse response functions for a 2 K temperature shock ($u_{1t}=2.0$).
-The plots illustrate the immediate response of the carbon cycle via the ICF and the long-term permanent shift in both variables.
-The secondary rise in temperature following the initial shock demonstrates the operation of the identified positive feedback loop between the thermal and carbon cycles.
-4.3.2 System Response to a 2 K Temperature Shock
-To evaluate the efficiency of the ICF and the robustness of the carbon cycle, we simulate a structural innovation in temperature of $u_{1t}=2.0$ K.
-This scenario, depicted in Figure 5, represents an abrupt warming event independent of the greenhouse effect, such as a shift in orbital forcing or ocean circulation.
-Upon the occurrence of the shock, the CO2 concentration immediately begins to rise, driven by the ICF coefficient $a_{21}=0.02082$.
-At the pre-industrial baseline of 280 ppm, this corresponds to an instantaneous release of approximately 5.8 ppm per degree of warming.
-In the initial 1,000-year step, the 2 K shock induces a rise of approximately 11.7 ppm in atmospheric CO2.
-The long-term trajectory shown in the left panels of Figure 5 reveals the operation of the cointegrating bond.
-As the temperature anomaly is sustained, the carbon cycle continues to adjust toward its new equilibrium level.
-Due to the property of cointegration in a non-stationary system, a temporary structural shock to temperature leads to a permanent shift in the CO2 baseline.
-The CO2 level eventually reaches a new steady state, representing the integrated response of ocean degassing and terrestrial biosphere adjustments.
-Crucially, the temperature itself exhibits a slight secondary increase after the initial shock.
-This is the structural representation of the positive feedback loop: the initial warming triggers a release of CO2 (ICF), which in turn amplifies the warming through the structural lag effect in subsequent millennia.
-The stabilization of the system at a higher baseline for both variables confirms that the identified VECM is dynamically stable and that the cointegrating relationship $T-18.748\ln C=\text{const}$ acts as a robust anchor for the Earth system’s long-term behavior.
+The immediate rise in CO2 (upper panels) illustrates the Contemporaneous Carbon Response (CCR), demonstrating that temperature immediately drives carbon fluctuations at the millennial scale.
 Concluding Remarks
-In this study, we have developed and estimated a Structural Vector Error Correction Model (SVECM) to disentangle the complex, bidirectional causalities between temperature and atmospheric CO2 concentrations during the Pleistocene.
-By focusing on the high-quality, gap-free record of Termination II and utilizing the strict exogeneity of orbital forcing for structural identification, we have provided a new perspective on the Earth system’s sensitivity and its response to carbon innovations.
-The primary finding of our analysis is the identification of a null Instantaneous Greenhouse Effect (IGE) at the millennial resolution.
-Our results consistently indicate that the contemporaneous parameter $a_{12}$ is statistically indistinguishable from zero.
-This does not imply that the greenhouse effect is physically absent; rather, it reveals that the radiative impact of CO2 is structurally localized as a delayed feedback.
-Our model identifies a powerful structural lag effect, $(\Gamma^{*}_{1})_{12}$, which delivers a warming of approximately 10.5 K in the millennium following a CO2 doubling.
-When combined with an equilibrium adjustment of approximately 1.3 K toward a long-run Earth System Sensitivity (ESS) of 13.0 K, the total millennial response reaches approximately 11.8 K.
-Crucially, the identification of $a_{12}=0$ is not a mere assumption but an empirical necessity derived from the data.
-The Instantaneous Carbon Feedback (ICF), estimated via first-difference OLS as $\hat{a}_{21}=0.02082$, lies at the precise center of the ICF Window for null IGE.
-The convergence between independent analytical paths—instrumental variable diagnostics for the carbon cycle and system-wide residual analysis for the thermal cycle—provides definitive empirical verification for the validity of our structural identification.
-While the median sensitivity from previous ensemble reconstructions such as Frank et al.
-(2010) sits at the boundary of statistical rejection for a null IGE, our data-driven estimate achieves maximal statistical consistency with the observed millennial dynamics.
-Furthermore, the robustness of our paleo-derived parameters is demonstrated by their remarkable predictive power for the modern era.
-The projected warming rate of 0.72 K per century, derived solely from our Termination II SVECM and adjusted for the modern CO2 accumulation factor of 0.61, is in near-perfect alignment with the instrumental observation of 0.73–0.80 K per century.
-The fact that a model calibrated on 130,000-year-old ice-core data can so accurately recover the scale of 21st-century climate change suggests that the structural properties of the Earth system’s coupling have remained fundamentally consistent across diverse climatic regimes.
-Our impulse response simulations further clarify the physical mechanism: the Earth system operates as a coupled engine where temperature changes, driven by orbital or other external shocks, immediately trigger a carbon response through the ICF.
-This released CO2 then acts as a powerful long-term regulator, driving the planet toward a new climatic baseline through delayed structural momentum and long-run equilibrium adjustment.
-In conclusion, this research bridges the gap between paleoclimatic structural analysis and modern climate monitoring.
-By identifying the specific temporal localization of the greenhouse effect, we provide a framework that explains both the millennial-scale phase-locking observed in the ice cores and the centennial-scale warming observed in the instrumental record.
-These findings underscore the critical importance of accounting for delayed feedbacks in long-term climate projections and demonstrate the enduring value of the Pleistocene record in deciphering the future of our planet.
+In this study, we have integrated rigorous time-series econometrics with paleoclimatic physics to re-examine the climate–carbon feedback architecture of the late Quaternary.
+By focusing on a gap-free 130,000-year interval and utilizing a structural VECM framework, we have circumvented the long-standing issues of interpolation bias and spurious regression that have often clouded the interpretation of ice-core records.
+Our primary finding is the identification of a fundamental asymmetry in the contemporaneous climate–carbon interaction at the millennial scale.
+The structural results demonstrate a unidirectional contemporaneous feedback: while temperature changes trigger an immediate and significant release of carbon from terrestrial and marine reservoirs (the CCR), the reciprocal radiative impact of CO2 on temperature (the CTR) is statistically negligible within the same 1,000-year time step.
+This discovery provides a novel bridge between conflicting paleoclimatic narratives.
+The apparent “synchronicity” observed in recent deglacial chronologies can be understood as the rapid manifestation of the carbon response (CCR), while the millennial-scale lag identified in earlier studies reflects the physical adjustment time required for the greenhouse effect to influence the Earth’s thermal state.
+Furthermore, our model identifies a stable Earth System Sensitivity (ESS) of 6.5 K per CO2 doubling.
+The fact that this estimate, derived through a dynamic structural framework, aligns so closely with previous OLS-based reconstructions suggests that the long-term coupling of the climate–carbon system is governed by a robust, persistent cointegrating trend that transcends simple glacial-interglacial oscillations.
+Finally, this millennial-scale perspective offers a sobering context for modern climate change.
+Our analysis shows that the 1.2 K global warming observed since the industrial revolution is not only consistent with the sensitivity scale found in paleo-data but also represents merely the first stage of a much larger, non-linear adjustment process.
+The discrepancy between current observations and our 1,000-year 5.9 K projection quantifies a massive “thermal debt” inherent in the Earth system.
+As the system moves toward the long-term equilibrium identified in this study, the commitment warming from centuries of cumulative slow feedbacks will likely present a challenge of unprecedented scale for the future of the planet.
 Appendix: Theoretical Background of VECM and Impulse Response
@@
 A critical implication of cointegration is the orthogonality between the permanent effect $\Xi$ and the cointegrating vector $\beta$:
-$\beta^{\prime}\Xi=0$ (20)
+$\beta^{\prime}\Xi=0$
 This mathematical property carries a profound physical meaning: while an exogenous shock (such as a sudden change in orbital forcing or a carbon cycle innovation) can permanently shift the absolute levels of temperature and CO2, it is fundamentally incapable of breaking the equilibrium "bond" between them.
@@
 Johansen (1991) provided the explicit representation for this permanent impact matrix:
-$\Xi=\beta_{\perp}(\alpha_{\perp}^{\prime}\Gamma\beta_{\perp})^{-1}\alpha_{\perp}^{\prime}$ (21)
+$\Xi=\beta_{\perp}(\alpha_{\perp}^{\prime}\Gamma\beta_{\perp})^{-1}\alpha_{\perp}^{\prime}$
 where $\Gamma=I-\sum_{i=1}^{p-1}\Gamma_{i}$.

### v4

@@
-arXiv:2601.06986v3[physics.ao-ph] 28 Feb 2026
+Structural Cointegration of the Climate–Carbon Feedback: Evidence from the Last 130,000 Years
 Abstract
-Using a gap-free, millennial-resolution record of atmospheric CO2 and Antarctic temperature since the Last Interglacial (130 ka ~), we structurally identify the feedback architecture of the climate–carbon system.
-The series are cointegrated, supporting estimation with a Vector Error Correction Model (VECM).
-The cointegrating relationship implies a temperature response of 13.0 K per CO2 doubling.
-Structural identification combining a Causality Allocation Curve (CAC) and Milankovitch-cycle instrumental variables yields a contemporaneous carbon response (CCR) of 5.77 ppm/K, while the contemporaneous temperature response (CTR) is statistically indistinguishable from zero.
+Using a gap-free, millennial-resolution ice-core record spanning the last 130,000 years, we identify the feedback architecture between Antarctic temperature and atmospheric CO2.
+The series are found to be cointegrated, justifying estimation with a Vector Error Correction Model (VECM).
+The estimated long-run relationship yields a temperature change of 13.0 K per CO2 doubling.
+Structural identification combining Milankovitch-cycle instrumental variables with the VECM residuals yields a contemporaneous carbon response (CCR) of 5.77 ppm/K, whereas the contemporaneous temperature response (CTR) is statistically indistinguishable from zero.
 Accounting for lagged feedbacks, the cumulative temperature response within one millennium of CO2 doubling reaches 11.8 K.
@@
 Establishing a robust causal architecture is not only vital for understanding past glacial-interglacial transitions but also for constraining long-term Earth System Sensitivity (ESS) in the context of anthropogenic warming.
-The statistical analysis of paleoclimate records faces three fundamental challenges.
-First, the high persistence of temperature and CO2 series carries the risk of spurious regression (Kaufmann et al., 2006; Cummins et al., 2022).
-Because these variables exhibit the properties of non-stationary, $I(1)$ processes, standard regression techniques can yield misleadingly significant correlations that lack physical causality.
-Second, a methodological schism exists regarding the appropriate framework for these series; while some advocate for a cointegration approach (Kaufmann et al., 2006; Miller, 2019), others argue that the system is ultimately mean-reverting and stationary, favoring standard Vector Autoregression (VAR) models (Davidson et al., 2016).
-Third, the irregular temporal spacing of ice-core proxies often necessitates linear interpolation, which has been shown to introduce substantial size distortions and biases in statistical testing (Miller, 2019).
-Beyond these statistical hurdles, a long-standing climate-physical debate persists regarding the “lead-lag” relationship between temperature and CO2.
-Early studies suggested that CO2 lagged Antarctic temperature by several centuries during deglaciations (Caillon et al., 2003), implying that carbon was a follower rather than a driver of initial warming.
-More recent re-evaluations, accounting for gas-ice age uncertainties, have proposed a nearly synchronous relationship (Parrenin et al., 2013).
-However, these findings often rely on reduced-form correlations that cannot isolate structural innovations from reciprocal feedbacks.
-While Lemoine (2010) utilized orbital parameters as instrumental variables (IV) to identify a significantly positive—yet relatively modest—climate–carbon feedback factor, the short-run structural dynamics within a unified millennial-scale framework remain under-explored.
-This study addresses these challenges by applying a Structural Vector Error Correction Model (SVECM) to a high-resolution, gap-free ice-core dataset.
-Our contribution is twofold.
-First, we identify a specific 130,000-year interval (the 130 ka sample) that is entirely free of observational gaps at a 1,000-year resolution.
-By utilizing this interval, we eliminate the need for interpolation, thereby physically neutralizing the statistical artifacts warned of by Miller (2019).
-Second, we employ a structural identification strategy that utilizes the strict exogeneity of Milankovitch cycles as an instrument to disentangle the contemporaneous temperature response (CTR, $a_{12}$) from the contemporaneous carbon response (CCR, $a_{21}$).
+The debate over the phase relationship between these variables has been long-standing.
+Caillon et al.
+(2003) famously reported that Antarctic temperature leads CO2 by approximately 800 years during Termination III, a finding often used to argue that the initial climate shift is driven by orbital forcing rather than greenhouse gases.
+However, as Lemoine (2010) pointed out, such simple lead-lag analyses based on peak-to-peak timing fail to disentangle the mutual causal effects from autocorrelation and reverse causality.
+In a complex system where temperature and carbon are mutually endogenous, a lead in one variable does not preclude the existence of a powerful, simultaneous feedback from the other.
+Without a framework that accounts for this reciprocity, the magnitude of the positive climate–carbon feedback may be significantly underestimated.
+Furthermore, the statistical analysis of paleoclimate records faces fundamental analytical challenges.
+A primary concern is the high persistence of temperature and CO2 series, which carries the risk of spurious regression (Kaufmann et al., 2006; Cummins et al., 2022).
+Because these variables exhibit non-stationary, $I(1)$ properties, a cointegration framework is essential to separate true physical bonds from stochastic trends.
+In this context, Kaufmann and Juselius (2013) provided a seminal benchmark by estimating a Cointegrated Vector Autoregression (CVAR) model; however, as noted by Miller (2019), such models often rely on interpolated data, which can introduce significant statistical artifacts.
+Furthermore, even when multivariate models are employed to capture slow internal dynamics (Kaufmann and Juselius, 2016), the identification of the contemporaneous interaction matrix ($A_{0}$) has historically relied on pre-assumed physical zero-constraints.
+Consequently, the objective allocation of causality within a single millennial time step—the very scale where the initial feedback triggers occur—remains an unresolved empirical question.
+In this paper, we address these challenges through two methodological advancements.
+First, we focus our primary structural analysis on a specific 130,000-year interval since the Last Interglacial that is entirely free of observational gaps at a 1,000-year resolution.
+By utilizing this gap-free sample, we eliminate the need for interpolation, thereby neutralizing the potential biases warned of by Miller (2019).
+Second, we employ a structural identification strategy that utilizes the strict exogeneity of Milankovitch cycles as an instrumental variable (IV).
+This allows us to statistically disentangle the contemporaneous temperature response (CTR, $a_{12}$) from the contemporaneous carbon response (CCR, $a_{21}$) without relying on pre-assumed physical constraints on the $A_{0}$ matrix.
 Our results reveal a stark asymmetry in the contemporaneous climate–carbon interaction.
-We find a robust, statistically significant CCR ($\hat{a}_{21}>0$), consistent with the magnitude suggested by last-millennium reconstructions (Frank et al., 2010).
+We find a robust, statistically significant CCR ($\hat{a}_{21}>0$), consistent with the magnitude suggested by last-millennium reconstructions (Frank et al., 2010) and the long-term sensitivity found by Kaufmann and Juselius (2013).
 In contrast, the CTR ($a_{12}$) is found to be statistically indistinguishable from zero at the millennial scale.
 This identification implies that while temperature changes trigger an immediate carbon release, the radiative feedback of CO2 onto temperature requires structural lags to manifest.
-This framework provides a unified explanation for the perceived “synchronicity” in deglacial records as a manifestation of the CCR, while reconciling the “lag” observed in other contexts as the physical adjustment time of the thermal system.
-By bridging the gap between time-series econometrics and paleoclimatic physics, we provide a more precise characterization of the feedbacks that govern the Earth’s long-term climate sensitivity.
+By bridging the gap between time-series econometrics and paleoclimatic physics, we provide a unified explanation for the perceived “synchronicity” in deglacial records as a manifestation of the CCR, while reconciling the “lag” observed in other contexts as the physical adjustment time of the thermal system.
 Model
@@
 In this study, we consider a three-variable system $y_{t}=(T_{t},\ln C_{t},M_{t})^{\prime}$, where $T_{t}$ is Antarctic temperature, $\ln C_{t}$ is the logarithm of CO2 concentration, and $M_{t}$ represents external orbital (Milankovitch) forcing.
-The VECM is expressed as:
-$\Delta y_{t}=\alpha\beta^{\prime}y_{t-1}+\Gamma_{1}\Delta y_{t-1}+\dots+\Gamma_{p-1}\Delta y_{t-p+1}+\epsilon_{t}$ (1)
-where $\beta$ contains the cointegrating vector, $\alpha$ represents the adjustment speeds to equilibrium, and $\Gamma_{i}$ captures the short-run lagged dynamics.
-The error term $\epsilon_{t}$ denotes the reduced-form innovations, which are generally correlated.
+We begin with a Structural Vector Autoregressive (SVAR) model that explicitly captures contemporaneous interactions ($A_{0}$):
+$y_{t}=A_{0}y_{t}+A_{1}y_{t-1}+\cdots+A_{p}y_{t-p}+u_{t},$ (1)
+where $u_{t}\sim\text{i.i.d.}~\mathcal{N}(0,\Sigma_{u})$.
+The corresponding reduced-form representation is given in Equation (6) in the Appendix.
+Comparing the two representations, we obtain $A_{i}=(I-A_{0})\Phi_{i}$ for $i=1,\dots,p$, and $u_{t}=(I-A_{0})\epsilon_{t}$.
+Moreover, the reduced-form VAR (6) can be reformulated as a VECM:
+$\Delta y_{t}=\alpha\beta^{\prime}y_{t-1}+\Gamma_{1}\Delta y_{t-1}+\dots+\Gamma_{p-1}\Delta y_{t-p+1}+\epsilon_{t},$ (2)
+where $\beta$ contains the cointegrating vector, $\alpha$ represents the adjustment speeds toward equilibrium, and $\Gamma_{i}$ captures the short-run lagged dynamics.
+The error term $\epsilon_{t}$ denotes the reduced-form innovations, which are generally correlated element-wise.
 Structural Identification and Causality Allocation
-To isolate structural shocks ($u_{t}$) from the reduced-form residuals ($\epsilon_{t}$), we define the relationship $(I-A_{0})\epsilon_{t}=u_{t}$, where $u_{t}$ is a vector of orthogonal structural innovations with $u_{t}\sim\text{i.i.d.}\mathcal{N}(0,\Sigma_{u})$.
+To isolate structural shocks ($u_{t}$) from the reduced-form residuals ($\epsilon_{t}$), we utilize the relationship $u_{t}=(I-A_{0})\epsilon_{t}$.
+This approach follows the framework established by Sims (1980), which provides a basis for analyzing endogeneity and structural innovations within a system of interrelated time series by utilizing the residuals of the estimated model.
 In our three-variable system, the contemporaneous interaction matrix $A_{0}$ is defined as:
@@
 $(1+a_{12}a_{21})s_{12}-a_{21}s_{11}-a_{12}s_{22}=0$ (3)
+where $s_{ij}$ denotes the sample covariance between $\epsilon_{it}$ and $\epsilon_{jt}$.
 Equation (3) demonstrates that the allocation of the total contemporaneous correlation between temperature and CO2 into directional causal links is uniquely determined by the sample covariance structure ($s_{ij}$) of the residuals.
@@
 The presence of cointegration ensures that temporary structural shocks lead to a permanent shift in the climatic baseline.
-The long-run impact matrix, defined as $\Xi=\lim_{h\to\infty}\Psi_{h}$ (see Appendix A, for details), provides a formal bridge between millennial-scale innovations and the long-term equilibrium states.
+The long-run impact matrix, defined as $\Xi=\lim_{h\to\infty}\Psi_{h}$ (see Appendix for details), provides a formal bridge between millennial-scale innovations and the long-term equilibrium states.
 Data
@@
 To further rigorously test the statistical significance of the CTR ($a_{12}$) while accounting for the uncertainty in the CCR ($a_{21}$), we perform a joint estimation using the Generalized Method of Moments (GMM).
-This approach incorporates both the regression model in Equation (3) and the orthogonality condition derived from the structural shock properties, known as the Causality Allocation Curve (CAC) or Equation (2).
+This approach incorporates both the regression model in Equation (4) and the orthogonality condition derived from the structural shock properties, i.e., the CAC, or Equation (3).
 The moment conditions for this joint estimation are defined by the structural identification framework established in Section 2.2.
@@
 The blue line represents the CAC derived from the residual covariance.
-The red point and dashed ellipse indicate the GMM joint estimate and the 95% confidence region.
+The red point and dashed ellipse indicate the GMM joint estimate (of Equations (3) and (4)) and the 95% confidence region.
 The intersection near the $a_{12}=0$ line confirms the unidirectional contemporaneous feedback from temperature to carbon.
@@
 First, to interpret the estimated reduced-form parameters ($\alpha,\beta,\Gamma$) in a structural context, we must consider the transformation $(I-A_{0})$.
-While the cointegrating vector $\beta$ and the adjustment speeds $\alpha$ remain invariant to this transformation due to the property $\beta^{\prime}\Xi=0$ (as formally derived in Appendix A), the short-run lag coefficients $\Gamma$ must be structurally re-identified using the contemporaneous parameters $a_{12}$ and $a_{21}$.
+While the cointegrating vector $\beta$ and the adjustment speeds $\alpha$ remain invariant to this transformation due to the property $\beta^{\prime}\Xi=0$ (as formally derived in Appendix), the short-run lag coefficients $\Gamma$ must be structurally re-identified using the contemporaneous parameters $a_{12}$ and $a_{21}$.
 By recalling the structural identity $(I-A_{0})\Delta y_{t}=(I-A_{0})\alpha\beta^{\prime}y_{t-1}+\Gamma^{*}_{1}\Delta y_{t-1}+\dots+u_{t}$, the relationship between the structural and reduced-form coefficients is defined by the identity $\Gamma^{*}_{1}=(I-A_{0})\Gamma_{1}$.
-The specific element for the CO22-to-temperature lag can be expanded as follows:
+The specific element for the $\ln C_{t-1}$ to $T_{t}$ effect can be expanded as follows:
 $(\Gamma^{*}_{1})_{12}=(\Gamma_{1})_{12}-a_{12}(\Gamma_{1})_{22}-a_{13}(\Gamma_{1})_{32}$ (5)
@@
 Consequently, the structural lag effect is directly identified by the estimated reduced-form coefficient: $(\Gamma^{*}_{1})_{12}=(\Gamma_{1})_{12}=15.201$.
-This structural consistency allows us to derive the Earth System Sensitivity (ESS) directly from the cointegrating relationship.
-The estimated long-term equilibrium for the 130 ka sample is given by $T-18.748\ln C=\text{const}$.
-This relationship implies that a doubling of atmospheric CO2 concentration is associated with a long-term temperature increase of $18.748\times\ln(2)\approx 13.0$ K in Antarctica.
-However, since this estimate is based on Antarctic ice core records, it reflects local temperature variations which are subject to polar amplification.
-Following the methodology of previous studies, we convert this local signal to a Global Average Temperature (GAT) equivalent by applying a scaling factor of approximately 0.5, as Antarctic temperature anomalies are known to be roughly twice as large as the global mean.
-This conversion yields a global ESS of approximately 6.5 K per CO2 doubling.
-This value is remarkably consistent with the reconstructed ESS of $9\pm 4.5$ K (95% CI) reported by Snyder (2016), confirming that our VECM-based structural identification captures the fundamental long-term coupling of the Earth’s climate and carbon cycle since the Last Interglacial.
-Beyond the long-term equilibrium, the structural VECM allows for a granular decomposition of the transient temperature response.
-We examine the impact of a doubling of atmospheric CO22 concentration ($\Delta\ln C=\ln(2)$) within a single millennial time step.
-Under the structural identification a12=0a{12}=0, the immediate temperature response after 1,000 years is driven by two distinct mechanisms: the structural lag effect and the error correction process.
-First, the structural lag effect $(\Gamma^{*}_{1})_{12}=15.201$ accounts for a temperature increase of $15.201\times\ln(2)\approx 10.5$ K.
-This represents the direct millennial-scale sensitivity of the climate system to a carbon shock.
-Second, the error correction term $\alpha=-0.103$ dictates the speed at which the system reverts to its long-term equilibrium.
-A doubling of CO2 creates a disequilibrium in the cointegrating relationship of $18.748\times\ln(2)\approx-13.0$ K.
-The resulting adjustment in the subsequent millennium contributes an additional warming of $(-0.103)\times(-13.0)\approx 1.3$ K.
-Summing these effects, the cumulative temperature response within one millennium of a CO2 doubling reaches $11.8$ K in Antarctica.
-In global average terms, this corresponds to approximately $5.9$ K of warming.
-While this magnitude appears large, it is remarkably consistent with the transient warming observed in the recent industrial era.
-For instance, the observed global temperature increase of approximately $1.2$ K over the last 150 years, associated with a CO2 rise from 280 to 428 ppm (a 53% increase), aligns with the sensitivity scale identified in our millennial-resolution model when accounting for the slower response times of the Earth system.
-Summing these effects, the cumulative temperature response within one millennium of a CO2 doubling reaches 11.8 K in Antarctica, which translates to a global average warming of approximately 5.9 K.
-While this 1,000-year response is substantial, it represents only the transient stage of the total adjustment toward the long-term equilibrium (ESS) of 6.5 K.
-The proximity of our VECM-based estimates to the results reported by Snyder (2016) warrants further methodological discussion.
-Snyder (2016) estimated the ESS using a simple OLS framework and was subsequently criticized for potentially ignoring the endogeneity of CO2 and the influence of orbital forcing.
-However, our identification of a stable cointegrating relationship ($\beta$) provides a strong theoretical justification for the robustness of such OLS-based reconstructions.
-In the presence of cointegration, OLS estimators possess the property of “super-consistency,” converging to the true long-run parameters even in the face of contemporaneous endogeneity.
-By employing a VECM, we have formally verified this long-term stability and separated the structural feedbacks from stochastic innovations.
-The fact that our structurally identified ESS (6.5 K) and millennial response (5.9 K) align closely with Snyder’s findings suggests that the climate–carbon coupling is so fundamentally dominated by the long-term cointegrating trend that the essence of the Earth system’s sensitivity is captured even by simpler models, provided they span sufficient glacial cycles.
-Our structural approach thus reinforces the empirical validity of previous reconstructions while providing a rigorous framework to account for internal dynamics and orbital exogeneity.
-This millennial sensitivity also provides a critical perspective on modern climate change.
-A global warming of 5.9 K per CO2 doubling after 1,000 years implies an average linear adjustment rate of approximately 0.885 K per 150 years.
-While this is numerically lower than the observed 1.2 K increase since the industrial revolution (associated with a 53% rise in CO2), this discrepancy is entirely consistent with the non-linear “front-loading” characteristic of climate responses.
-The immediate radiative forcing following a carbon shock triggers rapid warming in the upper ocean and atmosphere (often termed the Transient Climate Response), whereas our 6.5 K ESS estimate captures the slow, “committed” warming of the deep Earth system that unfolds over subsequent centuries.
-Rather than a contradiction, the gap between the modern 1.2 K response and our 1,000-year 5.9 K projection quantifies the massive, yet-to-be-realized thermal debt inherent in the Earth system’s long-term feedbacks.
-Finally, we evaluate our estimate of the Contemporaneous Carbon Response (CCR).
-Our GMM estimation yields $\hat{a}_{21}=0.0206$, which corresponds to a carbon cycle sensitivity of approximately $5.77$ ppm per 1 K of Antarctic warming.
-This result is highly consistent with the median estimate of $7.7$ ppm/K (likely range: 1.7–21.4 ppm/K) reported by Frank et al.
-(2010) for the last millennium.
-The stability of the CCR across different timescales—from the centennial resolution of the late Holocene to the millennial resolution of our 130 ka sample—underscores the robustness of the climate–carbon feedback architecture identified in this study.
+Our structural estimation yields an Antarctic Earth System Sensitivity (ESS) of 13.0 K, which is remarkably consistent with the 11.1 K reported by Kaufmann and Juselius (2013).
+This alignment is noteworthy given the methodological differences: while Kaufmann and Juselius (2013) utilized interpolated data and physical zero-constraints for identification, our results are derived from a gap-free sample with a data-driven GMM strategy.
+Specifically, the estimated long-term equilibrium for the 130 ka sample is given by $T-18.748\ln C=\text{const}$.
+This implies that a doubling of CO2 is associated with a long-term Antarctic increase of $18.748\times\ln(2)\approx 13.0$ K.
+Applying a polar amplification factor of approximately 0.5, this corresponds to a global ESS of $\sim$6.5 K, aligning with the reconstructed range of $9\pm 4.5$ K by Snyder (2016).
+Beyond the long-term equilibrium, the structural VECM reveals the specific causal architecture behind the “slow adjustment” of the climate system.
+While Kaufmann and Juselius (2016) identified adjustment scales of approximately 19 kyr, our identification of $a_{12}=0$ at the millennial scale explains this delay as a fundamental causal asymmetry.
+Under our null-CTR identification, the cumulative temperature response within one millennium of a CO2 doubling reaches 11.8 K in Antarctica (or 5.9 K globally).
+This response is driven by two mechanisms: the structural lag effect $(\Gamma^{*}_{1})_{12}\times\ln(2)\approx 10.5$ K and the error correction process $\alpha\times(-13.0)\approx 1.3$ K.
+This result implies that while temperature triggers an immediate carbon release (CCR), the reciprocal radiative impact of CO2 onto temperature is not instantaneous but manifests through structural lags.
+Finally, we compare these paleo-derived responses with modern climate observations.
+A 1,000-year response of 5.9 K global warming may appear high compared to the 1.2 K increase observed since the industrial revolution.
+However, this discrepancy is consistent with the non-linear “front-loading” of climate response: the immediate warming (Transient Climate Response) is followed by centuries of “committed” warming from slow internal feedbacks.
+The gap between the modern 1.2 K and our 1,000-year projection quantifies the massive “thermal debt” inherent in the Earth system’s long-term dynamics.
+Our GMM estimate $\hat{a}_{21}=0.0206$ corresponds to a carbon cycle sensitivity of $\sim$5.77 ppm/K, which is highly consistent with the $7.7$ ppm/K reported by Frank et al.
+(2010).
+The stability of this CCR across different frameworks reinforces the warning by Kaufmann and Juselius (2016) that univariate analyses are prone to bias.
+By explicitly modeling the joint endogeneity, we ensure that our sensitivity parameters reflect the integrated dynamics of the climate–carbon system rather than isolated correlations.
 Impulse Response Simulation of Climate–Carbon Dynamics
@@
 Figure 4 illustrates the response to the CO2 doubling shock.
-Consistent with our null-CTR identification ($a_{12}=0$), temperature shows no immediate response within the same 1,000-year step.
+Consistent with our null-CTR identification, temperature shows no immediate response within the same 1,000-year step.
 Instead, the warming is triggered by structural lag effects, gradually converging toward the long-term ESS of 13.0 K in Antarctica.
@@
 As the system moves toward the long-term equilibrium identified in this study, the commitment warming from centuries of cumulative slow feedbacks will likely present a challenge of unprecedented scale for the future of the planet.
-Appendix: Theoretical Background of VECM and Impulse Response
+Appendix: The VECM Representation and Impulse Responses
 Decay of Shocks in a Stationary VAR
@@
 Assuming a constant shock has been applied from the past to the present ($\epsilon_{t-h}=\epsilon$), Equation (7) yields $y_{t}=(\sum_{h=0}^{\infty}\Psi_{h})\epsilon$.
-On the other hand, assuming a long-run equilibrium state ($y_{t}=y_{t-1}=\dots=y$) and substituting this into Equation (5), we obtain $\left[I-\sum_{i=1}^{p}\Phi_{i}\right]y=Gy=\epsilon$.
+On the other hand, assuming a long-run equilibrium state ($y_{t}=y_{t-1}=\dots=y$) and substituting this into Equation (6), we obtain $\left(I-\sum_{i=1}^{p}\Phi_{i}\right)y=Gy=\epsilon$.
 By comparing the two, the following relationship is derived:
@@
 $\Delta y_{t}=C_{0}\epsilon_{t}+C_{1}\epsilon_{t-1}+C_{2}\epsilon_{t-2}+\dots=\sum_{j=0}^{\infty}C_{j}\epsilon_{t-j}\quad(C_{0}=I)$ (9)
-where $C_{j}=\Psi_{j}-\Psi_{j-1}$ represents the impact of the shock $\epsilon_{t-j}$ on the current "change" $\Delta y_{t}$.
+where $C_{j}=\Psi_{j}-\Psi_{j-1}$ represents the impact of the shock $\epsilon_{t-j}$ on the current “change” $\Delta y_{t}$.
 Given that the level $y_{t}$ is the cumulative sum of past differences $y_{t}=y_{0}+\sum_{j=1}^{t}\Delta y_{j}$, the impulse response for the level, $\Psi_{h}=C_{h}+\Psi_{h-1}$, can be rewritten as:
@@
 The impulse response after $h$ periods can be decomposed as:
-$\Psi_{h}=\sum_{j=0}^{h}C_{j}=\left(\sum_{j=0}^{\infty}C_{j}\right)-\left(\sum_{j=h+1}^{\infty}C_{j}\right)=\Xi+\Upsilon_{h}$
+$\Psi_{h}=\sum_{j=0}^{h}C_{j}=\sum_{j=0}^{\infty}C_{j}-\sum_{j=h+1}^{\infty}C_{j}=\Xi+\Upsilon_{h}$
 From Equation (10), it is evident that $\lim_{h\to\infty}\Upsilon_{h}=\lim_{h\to\infty}(\Psi_{h}-\Xi)=0$, meaning $\Upsilon_{h}$ represents the decaying transitory component.
@@
 $\Delta y_{t}=\Pi y_{t-1}+\Gamma_{1}\Delta y_{t-1}+\dots+\Gamma_{p-1}\Delta y_{t-p+1}+\epsilon_{t}$ (11)
-where $\Pi=\alpha\beta^{\prime}$ represents the long-run equilibrium (cointegrating) relationships.
 Note that $\Phi_{1}=I+\Pi+\Gamma_{1}$, $\Phi_{i}=\Gamma_{i}-\Gamma_{i-1}$, and $\Phi_{p}=-\Gamma_{p-1}$.
-Engle and Granger (1987) demonstrated that for $I(1)$ series, a linear combination $\beta^{\prime}y_{t}$ exists that is stationary ($I(0)$).
+Given the non-stationarity of the level series $y_{t-1}$, $\Pi$ cannot be of full-rank.
+The rank-dificient case $\Pi=\alpha\beta^{\prime}$ represents the long-run equilibrium (cointegrating) relationships.
+Engle and Granger (1987) demonstrated that for an $I(1)$ series $y_{t}$, a linear combination $\beta^{\prime}y_{t}$ exists that is stationary, since $\Delta y_{t}$ in Equation (11) must all be $I(0)$.
 A critical implication of cointegration is the orthogonality between the permanent effect $\Xi$ and the cointegrating vector $\beta$:
 $\beta^{\prime}\Xi=0$
-This mathematical property carries a profound physical meaning: while an exogenous shock (such as a sudden change in orbital forcing or a carbon cycle innovation) can permanently shift the absolute levels of temperature and CO2, it is fundamentally incapable of breaking the equilibrium "bond" between them.
+This property carries a profound physical meaning: while an exogenous shock (such as a sudden change in orbital forcing or a carbon cycle innovation) can permanently shift the absolute levels of temperature and CO2, it is fundamentally incapable of breaking the equilibrium “bond” between them.
 The variables move to a new climatic baseline, but the underlying physical relationship, defined by $\beta$, remains intact.